### v2

Abstract
The Arnoldi process provides an efficient framework for approximating functions of a matrix applied to a vector, i.e., of the form $f(M)\bm{b}$, by repeated matrix-vector multiplications. In this paper, we derive an a priori error estimate for approximating the action of a matrix square root using the Arnoldi process, where the integral representation of the error is reformulated in terms of the error for solving the linear system $M\bm{x}=\bm{b}$. The results extend the error analysis of the Lanczos method for Hermitian matrices in [Chen et al., SIAM J. Matrix Anal. Appl., 2022] to non-Hermitian cases. Furthermore, to make the method applicable to large-scale problems, we assume that the matrices are preprocessed utilizing data-sparse approximations preserving positive definiteness, and then establish a refined error bound in this setting. The numerical results on matrices with different structures demonstrate that our theoretical analysis yields a reliable upper bound. Finally, simulations on large-scale matrices arising in particulate suspensions validate the effectiveness and practicality of the approach.
keywords:
Arnoldi approximation, matrix square root, matrix functions
{MSCcodes}
65F60 65Z05 70-10
1 Introduction
Given a matrix $M\in\mathbb{C}^{n\times n}$, its square roots are the solutions of $X^{2}=M.$ If $M$ is positive definite, then $M$ admits a unique positive definite square root [19], denoted as $M^{1/2}$. Matrix square roots frequently arise in stochastic simulation [12], state estimation [2], image reconstruction [3], and Bayesian optimization [31, 18]. Hence, the computation of a matrix square root is a fundamental problem in numerical linear algebra.
There are a list of approaches in the literature for computing the square root of a matrix. Cholesky decomposition and eigenvalue decomposition provide useful tools for symmetric positive definite (SPD) matrices [35]. For general matrices, the Schur decomposition transforms the matrix into a triangular form, where the square root can be computed explicitly [19]. Additionally, Newton-like iterations [19] provide stable and efficient implementation for approximating the matrix square root. Unfortunately, these classical methods all suffer from at least $\mathcal{O}(n^{3})$ complexity for large-scale matrices [19]. On the other hand, even when $M$ is sparse or structured, its square root is typically dense, making direct computation and storage costly or even infeasible for large-scale problems. Consequently, rather than computing the square root of a matrix explicitly, the main interest lies in computing the action of a matrix square root on a vector. This is also the case for general matrix functions. Krylov subspace methods such as the Arnoldi iteration, therefore provide a powerful framework, as they only rely on repeated matrix-vector multiplications to construct the Krylov subspace. For general dense matrices, the overall computational cost of the Arnoldi process is about $\mathcal{O}(kn^{2})$, where $k$ is the number of Arnoldi iterations. When the matrix is Hermitian, the Arnoldi process simplifies to the Lanczos algorithm, characterized by a three-term recurrence. Strategies such as restarting [13, 15, 20], truncated orthogonalization [30, 9], and randomized sketching [30, 9, 29, 16] have been explored to reduce memory and orthogonalization cost.
Recently, there has been growing interest in the theoretical analysis of Krylov methods for approximating the action of matrix functions. As shown in [6], the error of the Lanczos algorithm after $k$ iterations for approximating $f(M)\mathbf{b}$ is bounded by the error of the best $k$-degree polynomial that uniformly approximates $f(x)$ over the spectral interval $[\lambda_{\min}(M),\lambda_{\max}(M)]$, where $\lambda_{\min}(M)$ and $\lambda_{\max}(M)$ denote the smallest and largest eigenvalues of the matrix $M$, respectively. Such a classical bound has also been studied within the setting of finite precision arithmetic [28]. In [11], a posteriori error estimates for polynomial Krylov methods are developed and applied to a broad class of matrix functions with integral representations. Based on the Cauchy integral formula, Chen et al. [6] establish a priori and a posteriori error bounds for the Lanczos method by reformulating the approximation error in terms of the error of solving a shifted linear system. In addition, the authors in [1] explore the near optimality of the Lanczos methods for rational functions with real poles lying outside the spectral interval of $M$. Their results are established under the assumption that the matrices are Hermitian, while in many practical applications, the matrices of interest are non-Hermitian. This motivates us to extend their results to the non-Hermitian setting.
In this work, we present an a priori error bound of the Arnoldi process specifically for evaluating a matrix square root, extending the results in [6] to the case of positive-definite non-Hermitian matrices. The main idea is to utilize the integral representation of the error in the iterations of the Arnoldi process, and then relate it to the error for solving the linear system $M\bm{x}=\bm{b}$. Compared to Hermitian matrices, non-Hermitian matrices exhibit complex spectral behavior: they are not necessarily diagonalizable, and even when they are, the eigenvectors are typically not orthogonal and the eigenvalues may be complex. To address these difficulties, we employ two techniques: the first is applying the Cauchy integral formula to estimate the norms of matrix functions for general matrices that may not be diagonalizable; the second is relating the modulus of a function in the complex plane to the real-valued function of the modulus, reducing the analysis to the positive real line. Moreover, for large-scale problems, we derive an error bound for approximating the matrix square root based on a data-sparse representation of the matrix. In numerical experiments, we investigate matrices with different structures to verify the theoretical results. Furthermore, we validate the effectiveness of our approach on large-scale mobility matrices arising from particulate suspension simulations. To improve computational efficiency and reduce storage cost, the matrices are preprocessed through adaptive cross approximation and stored in the form of hierarchical matrices.
The remainder of this paper is arranged as follows. In Section 2, we introduce the Arnoldi process for matrix functions. Error estimation of the Arnoldi process for computing a square root is provided in Section 3. Section 4 presents the implementation and reports our numerical results. Finally, concluding remarks are made in Section 5.
2 Matrix functions and Arnoldi iterations
There are different ways to define a function of a matrix, depending on the properties of the matrix and the type of function. In this work, we focus on two of them: the definition via the Jordan canonical form and the Cauchy integral representation.
Given a square matrix $M\in\mathbb{C}^{n\times n}$, assume its Jordan canonical form is given by
$M=ZJZ^{-1},$
where $Z$ is a nonsingular matrix, and $J=\operatorname{diag}\left(J_{1},J_{2},\ldots,\right.\left.J_{p}\right)$ with $J_{k}\in\mathbb{R}^{m_{k}}$, $\sum_{i=1}^{p}m_{k}=n$ being a Jordan block associated with the eigenvalue $\lambda_{k}$, $1\leq k\leq p$. Then, the matrix function $f(M)$ is defined as
$f(M):=Zf(J)Z^{-1}=Z\operatorname{diag}\left(f\left(J_{1}\right),f\left(J_{2}\right),\cdots,f\left(J_{p}\right)\right)Z^{-1},$
where $f(J_{k})$ takes the form:
(5) $f\left(J_{k}\right):=\left[\begin{array}[]{cccc}f\left(\lambda_{k}\right)&f^{\prime}\left(\lambda_{k}\right)&\ldots&\frac{f^{\left(m_{k}-1\right)}\left(\lambda_{k}\right)}{\left(m_{k}-1\right)!}\\ &f\left(\lambda_{k}\right)&\ddots&\vdots\\ &&\ddots&f^{\prime}\left(\lambda_{k}\right)\\ &&&f\left(\lambda_{k}\right)\end{array}\right].$
Alternatively, when $f$ is analytic on a domain containing the spectrum of $M$, one can also define $f(M)$ using the Cauchy integral formula:
$f(M)=-\frac{1}{2\pi i}\int_{\mathcal{C}}f(z)(M-zI)^{-1}\,\mathrm{d}z,$
where $\mathcal{C}$ is a closed contour enclosing the spectrum of $M$. The two definitions are equivalent when $f$ is analytic [19].
Algorithm 1 Arnoldi Process
1:  Input: Matrix $M\in\mathbb{C}^{n\times n}$, vector $\bm{b}\in\mathbb{C}^{n}$, and iteration number $k$
2:  Initialize: Set $\bm{q}_{1}=\bm{b}/\|\bm{b}\|$ and $Q_{k}=[\bm{q}_{1}]$
3:  for $j=1$ to $k$ do
4:     $\bm{w}=M\bm{q}_{j}$
5:     for $i=1,\dots,j$ do
6:        $h_{i,j}=\bm{q}_{i}^{*}\bm{w}$
7:        $\bm{w}=\bm{w}-h_{i,j}\bm{q}_{i}$
8:     end for
9:     $h_{j+1,j}=\|\bm{w}\|$
10:     if $h_{j+1,j}=0$ then
11:        Stop
12:     end if
13:     $\bm{q}_{j+1}=\bm{w}/h_{j+1,j}$
14:     Append $\bm{q}_{j+1}$ to $Q_{k}=[\bm{q}_{1},\dots,\bm{q}_{j+1}]$
15:  end for
In practice, we are primarily interested in the action of a matrix function $f(M)$ on a vector $\bm{b}$. To achieve this efficiently, we employ the Arnoldi process, which is described in Algorithm 1. Given $\bm{b}\in\mathbb{C}^{n}$, this yields
(6) $MQ_{k}=Q_{k}H_{k}+h_{k+1,k}\bm{q}_{k+1}\bm{e}_{k}^{T},$
where $H_{k}=[h_{i,j}]\in\mathbb{C}^{k\times k}$ is an upper Hessenberg matrix satisfying $Q_{k}^{*}MQ_{k}=H_{k}$, $\bm{e}_{k}$ is the $k$-th canonical basis vector, and $Q_{k}$ is an orthonormal matrix whose columns, $\bm{q}_{i}$, form a basis for the $k$-th Krylov space,
$\mathcal{K}_{k}(M,\bm{b})=\text{span}\{\bm{b},M\bm{b},M^{2}\bm{b},\dots,M^{k-1}\bm{b}\}.$
Then, the $k$-th Arnoldi iteration for approximating $f(M)\bm{b}$ is given by
$f(M)\bm{b}\approx\operatorname{Arn}_{k}(f;M,\bm{b}):=Q_{k}f(H_{k})Q_{k}^{*}\bm{b}=\|\bm{b}\|Q_{k}f(H_{k})\bm{e}_{1}.$
Note that when $f(x)=1/x$, the Arnoldi approximation reduces to the Full Orthogonalization Method (FOM) [33] for solving the linear system $M\bm{x}=\bm{b}$, with initial guess $\bm{x}_{0}=\bm{0}$. For the rest of this paper, whenever FOM is referred, we always assume the initial guess is zero.
3 Error bound
In this section, we study an a priori error bound of the Arnoldi process for approximating general matrix functions and then focus on the analysis of the matrix square root.
3.1 Integral representation of the error
Let $f$ be analytic along and inside a contour $\mathcal{C}$ that encloses the spectrum of both $M$ and $H_{k}$. By the Cauchy integral representation of $f(H_{k})$, the $k$-th Arnoldi approximation to $f(M)\bm{b}$ is given by:
$\operatorname{Arn}_{k}(f;M,\bm{b})=\|\bm{b}\|Q_{k}f(H_{k})\bm{e}_{1}=-\frac{\|\bm{b}\|}{2\pi i}\int_{\mathcal{C}}f(z)Q_{k}(H_{k}-zI)^{-1}\bm{e}_{1}\,\mathrm{d}z.$
The integral involves solving a shifted linear system $(M-zI)\bm{x}=\bm{b}$ using FOM, yielding the approximation $\|\bm{b}\|Q_{k}(H_{k}-zI)^{-1}\bm{e}_{1}$. The corresponding error and residual are defined as:
$\bm{\xi}_{z}^{k}:=(M-zI)^{-1}\bm{b}-Q_{k}(H_{k}-zI)^{-1}\|\bm{b}\|\bm{e}_{1}\quad\text{and}\quad\bm{r}_{z}^{k}:=(M-zI)\bm{\xi}_{z}^{k}.$
Then, we represent the error of the Arnoldi process for $f(M)\bm{b}$ by:
(7) $f(M)\bm{b}-\operatorname{Arn}_{k}(f;M,\bm{b})=f(M)\bm{b}-\|\bm{b}\|Q_{k}f(H_{k})\bm{e}_{1}=-\frac{1}{2\pi i}\int_{\mathcal{C}}f(z)\bm{\xi}_{z}^{k}\,\mathrm{d}z.$
Lemma 3.1.
For $z\in\mathbb{C}$ lying on a closed contour $\mathcal{C}$ that encloses the spectrum of $H_{k}$,
$\bm{r}_{z}^{k}=\left(\frac{(-1)^{k}}{\det(H_{k}-zI)}\prod_{j=1}^{k}h_{j+1,j}\right)\|\bm{b}\|\,\bm{q}_{k+1}.$
Proof 3.2.
According to Algorithm 1, we have
$(M-zI)Q_{k}=Q_{k}(H_{k}-zI)+h_{k+1,k}\bm{q}_{k+1}\bm{e}_{k}^{T},$
and then,
$\|\bm{b}\|(M-zI)Q_{k}(H_{k}-zI)^{-1}\bm{e}_{1}$ $=Q_{k}\|\bm{b}\|\bm{e}_{1}+h_{k+1,k}\bm{q}_{k+1}\bm{e}_{k}^{T}(H_{k}-zI)^{-1}\|\bm{b}\|\bm{e}_{1}$
$=\bm{b}+\|\bm{b}\|h_{k+1,k}\bm{q}_{k+1}\bm{e}_{k}^{T}(H_{k}-zI)^{-1}\bm{e}_{1}.$
Noting that $(H_{k}-zI)^{-1}=(1/\det(H_{k}-zI))\text{adj}(H_{k}-zI)$ and using the properties of an adjugate matrix leads to
$\bm{e}_{k}^{T}(H_{k}-zI)^{-1}\bm{e}_{1}=\frac{\text{cof}(H_{k}-zI)_{1k}}{\det(H_{k}-zI)}=\frac{(-1)^{k+1}}{\det(H_{k}-zI)}\prod_{j=1}^{k-1}h_{j+1,j},$
where $\text{cof}(H_{k}-zI)_{1k}$ denotes the $(1,k)$-cofactor of $H_{k}-zI$. Substituting this into the expression for $\bm{r}_{z}^{k}$ gives the desired result.
Applying Lemma 3.1, we establish the following result, which connects the errors and residuals of the Arnoldi approximations for different shifts.
Lemma 3.3.
Assuming $M$ and $M-zI$ are invertible for some $z\in\mathbb{C}$, we obtain
$\bm{\xi}_{z}^{k}=\det(h_{z}(H_{k}))h_{z}(M)\bm{\xi}_{0}^{k}\quad\text{and}\quad\bm{r}_{z}^{k}=\det(h_{z}(H_{k}))\bm{r}_{0}^{k},$
where the function $h_{z}(x)$ is defined as $h_{z}(x):=(x-z)^{-1}x$.
Proof 3.4.
The proof follows a similar argument to that of Corollary 2.4 in [6].
Then, we have the following Theorem.
Theorem 3.5.
Given a nonsingular matrix $M\in\mathbb{C}^{n\times n}$ and a vector $\bm{b}\in\mathbb{C}^{n}$, let function $f$ be analytic on and inside a closed contour $\mathcal{C}$ enclosing the spectrum of $M$ and $H_{k}$, where $H_{k}$ is an upper Hessenberg matrix generated from the Arnoldi process for $M$ and $\bm{b}$. Then, the error of the Arnoldi iteration for $f(M)b$ can be expressed as
(8) $f(M)\bm{b}-\operatorname{Arn}_{k}(f;M,\bm{b})=\left(-\frac{1}{2\pi i}\int_{\mathcal{C}}f(z)\operatorname{det}\left(h_{z}\left(H_{k}\right)\right)h_{z}(M)\,\mathrm{d}z\right)\bm{\xi}_{0}^{k}.$
Proof 3.6.
The result is immediate from Eq. 7 and Lemma 3.3:
$f(M)\bm{b}-\operatorname{Arn}_{k}(f;M,\bm{b})$ $=-\frac{1}{2\pi i}\int_{\mathcal{C}}f(z)\bm{\xi}_{z}^{k}\,\mathrm{d}z$
$=-\frac{1}{2\pi i}\int_{\mathcal{C}}f(z)\det(h_{z}(H_{k}))h_{z}(M)\bm{\xi}_{0}^{k}\,\mathrm{d}z.$
Remark 3.7.
Theorem 3.5 shows that for any nonsingular $M$, the error of the Arnoldi iteration for computing $f(M)\bm{b}$ is reformulated as the error in solving the linear system $M\bm{x}=\bm{b}$. This further simplifies the contour integral. For a more general setting, the error admits a bound related to the error of solving shifted linear systems of the form $(M-wI)\bm{x}=\bm{b}$, where $w$ lies outside the spectrum of both $H_{k}$ and $M$; see Corollary 2.5 in [6] for details.
3.2 Matrix square root
We now turn to deriving an a priori error estimate of the Arnoldi process for approximating a matrix square root, that is $f(x)=x^{1/2}$. To ensure the existence of a unique positive definite square root, we assume that $M$ is positive definite. Here, the matrices of interest are non-Hermitian. For such matrices, the eigenvalues may lie off the real axis leading to a contour integral in the complex plane for the error of the Arnoldi process, thereby introducing challenges. We derive the error bound starting from Theorem 3.5, and organize the proof into the following steps:
Step 1: Select an appropriate contour in the complex plane, and derive an initial error bound for the 2-norm of the Arnoldi approximation.
Step 2: Bound the modulus of the complex function $h_{z}(w)$ using a real-valued function of the modulus, and then estimate $\|h_{z}(M)\|$ with a new contour integral.
Step 3: Apply the properties of the Beta and Gamma functions to obtain an error bound in terms of the residual $\bm{r}_{0}^{k}$ for solving $Mx=b$.
Step 4: Conclude the proof by deriving a upper bound for $\bm{r}_{0}^{k}$.
From this, we establish the following theorem.
Theorem 3.8.
Given a positive-definite matrix $M\in\mathbb{C}^{n\times n}$ and a vector $\bm{b}\in\mathbb{C}^{n}$, the relative error of the Arnoldi process for approximating the matrix square root is bounded by
$\frac{\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};M,\bm{b})\|}{\|M^{1/2}\bm{b}\|}\leq\frac{2\sqrt{2}(\kappa(M))^{5/2}}{(k-1/2)^{3/4}}\frac{\|\bm{r}^{k}_{0}\|}{\|\bm{b}\|}\leq\frac{4\sqrt{2}(\kappa(M))^{7/2}}{(k-1/2)^{3/4}},$
where $\kappa(M)$ denotes the condition number of $M$, defined as $\kappa(M)=\frac{\sigma_{\max}(M)}{\sigma_{\min}(M)}$, and $\sigma_{\max}(M)$ and $\sigma_{\min}(M)$ denote the largest and smallest singular values of $M$, respectively
Proof 3.9.
Step 1: Choose an Appropriate Contour. Define $\mathcal{C}$ as a keyhole contour, consisting of a large circular arc $\mathcal{C}_{R}$ of radius $R$ centered at the origin, a small indented circular arc $\mathcal{C}_{\epsilon}$ of radius $\epsilon$ around the origin, and two horizontal line segments $\mathcal{C}^{+}_{\delta}$ and $\mathcal{C}^{-}_{\delta}$, which run at a small vertical distance $0<\delta<\epsilon$ above and below the branch cut, respectively (see Fig. 1 for details).
[…]
Figure 1: Keyhole contour with outer circle $\mathcal{C}_{R}$, inner circle $\mathcal{C}_{\epsilon}$, and offset $\delta$ above/below the branch cut.
Taking the 2-norm on both sides of (8) with $f(z)=z^{1/2}$, we obtain
$\left\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(f;M,\bm{b})\right\|\leq$ $\left(\frac{1}{2\pi}\int_{\mathcal{C}_{R}}|z^{1/2}|\cdot\left|\operatorname{det}\left(h_{z}\left(H_{k}\right)\right)\right|\cdot\left\|h_{z}(M)\right\|\,|\mathrm{d}z|\right)\left\|\bm{\xi}_{0}^{k}\right\|$
$-\left(\frac{1}{2\pi}\int_{\mathcal{C}_{\epsilon}}|z^{1/2}|\cdot\left|\operatorname{det}\left(h_{z}\left(H_{k}\right)\right)\right|\cdot\left\|h_{z}(M)\right\|\,|\mathrm{d}z|\right)\left\|\bm{\xi}_{0}^{k}\right\|$
$+\left(\frac{1}{2\pi}\int_{\mathcal{C}_{\delta}^{+}}|z^{1/2}|\cdot\left|\operatorname{det}\left(h_{z}\left(H_{k}\right)\right)\right|\cdot\left\|h_{z}(M)\right\|\,|\mathrm{d}z|\right)\left\|\bm{\xi}_{0}^{k}\right\|$
(9) $-\left(\frac{1}{2\pi}\int_{\mathcal{C}_{\delta}^{-}}|z^{1/2}|\cdot\left|\operatorname{det}\left(h_{z}\left(H_{k}\right)\right)\right|\cdot\left\|h_{z}(M)\right\|\,|\mathrm{d}z|\right)\left\|\bm{\xi}_{0}^{k}\right\|.$
Let $\mathcal{I}_{1},\ldots,\mathcal{I}_{k}\subset\mathbb{C}$ be subsets such that the eigenvalues of $H_{k}$ are contained in $\mathcal{I}_{i}$ for $i=1,\ldots,k$. Define the norm $\|h\|_{\mathcal{I}}:=\sup_{x\in\mathcal{I}}|h(x)|$ for a function $h(x)$ on $\mathcal{I}\subset\mathbb{C}$, where $|h(x)|$ denotes the modulus of $h(x)$. Note that the Jordan block definition of matrix functions Eq. 5 implies the eigenvalues of $f(M)$ are exactly $f(\lambda_{i})$, where $\lambda_{i}$ are the eigenvalues of $M$. Thus,
$\left|\operatorname{det}\left(h_{z}\left(H_{k}\right)\right)\right|=\left|\prod_{i=1}^{k}\lambda_{i}\left(h_{z}\left(H_{k}\right)\right)\right|=\left|\prod_{i=1}^{k}h_{z}\left(\lambda_{i}\left(H_{k}\right)\right)\right|\leq\prod_{i=1}^{k}\left\|h_{z}\right\|_{\mathcal{I}_{i}}.$
Plugging this into (3.9) yields
$\left\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(f;M,\bm{b})\right\|\leq$ $\left(\frac{1}{2\pi}\int_{\mathcal{C}_{R}}|z^{1/2}|\cdot\prod_{i=1}^{k}\left\|h_{z}\right\|_{\mathcal{I}_{i}}\cdot\left\|h_{z}(M)\right\|\,|\mathrm{d}z|\right)\left\|\bm{\xi}_{0}^{k}\right\|$
$-\left(\frac{1}{2\pi}\int_{\mathcal{C}_{\epsilon}}|z^{1/2}|\cdot\prod_{i=1}^{k}\left\|h_{z}\right\|_{\mathcal{I}_{i}}\cdot\left\|h_{z}(M)\right\|\,|\mathrm{d}z|\right)\left\|\bm{\xi}_{0}^{k}\right\|$
$+\left(\frac{1}{2\pi}\int_{\mathcal{C}_{\delta}^{+}}|z^{1/2}|\cdot\prod_{i=1}^{k}\left\|h_{z}\right\|_{\mathcal{I}_{i}}\cdot\left\|h_{z}(M)\right\|\,|\mathrm{d}z|\right)\left\|\bm{\xi}_{0}^{k}\right\|$
(10) $-\left(\frac{1}{2\pi}\int_{\mathcal{C}_{\delta}^{-}}|z^{1/2}|\cdot\prod_{i=1}^{k}\left\|h_{z}\right\|_{\mathcal{I}_{i}}\cdot\left\|h_{z}(M)\right\|\,|\mathrm{d}z|\right)\left\|\bm{\xi}_{0}^{k}\right\|.$
As the outer radius $R\to\infty$, the integral over the large circular arc vanishes for $k\geq 1$ since $|z^{1/2}|=O\left(R^{1/2}\right)$, $\left\|h_{z}\right\|_{\mathcal{I}_{i}}=O\left(R^{-1}\right)$, and $\|h_{z}(M)\|=O(R^{-1})$. Similarly, as the inner radius $\epsilon\to 0$, the integral goes to zero. Thus, as $\epsilon\to 0$ and $R\to\infty$, the contributions to the integral come from $\mathcal{C}^{+}_{\delta}$ and $\mathcal{C}^{-}_{\delta}$. Also, as $\epsilon\to 0$, $\delta$ approaches 0. Then, we have
$\left\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(f;M,\bm{b})\right\|$
$\leq$ $\lim_{\begin{subarray}{c}\epsilon\to 0\\ R\to\infty\end{subarray}}\left(\frac{1}{2\pi}\int_{-R}^{-\epsilon}\left|(x+\delta i)^{1/2}\right|\cdot\left(\prod_{i=1}^{k}\left\|h_{x+\delta i}\right\|_{\mathcal{I}_{i}}\right)\cdot\left\|h_{x+\delta i}(H_{k})\right\|\mathrm{~{}d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|$
$-\left(\frac{1}{2\pi}\int^{-R}_{-\epsilon}\left|(x-\delta i)^{1/2}\right|\cdot\left(\prod_{i=1}^{k}\left\|h_{x-\delta i}\right\|_{\mathcal{I}_{i}}\right)\cdot\left\|h_{x-\delta i}(H_{k})\right\|\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|$
$=$ $\left(\frac{1}{\pi}\int_{0}^{\infty}x^{1/2}\cdot\left(\prod_{i=1}^{k}\left\|h_{-x}\right\|_{\mathcal{I}_{i}}\right)\cdot\left\|h_{-x}(M)\right\|\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|.$
Step 2: Bound $\left\|h_{-x}\right\|_{\mathcal{I}_{i}}$ and $\left\|h_{-x}(M)\right\|$. Consider any $w\in\mathbb{C}$ and a fixed $x\in\mathbb{R}$ such that $x>0$, then,
$|w+x|=\sqrt{(\operatorname{Re}(w)+x)^{2}+\operatorname{Im}(w)^{2}}\quad\text{and}\quad\sqrt{|w|^{2}+x^{2}}=\sqrt{\operatorname{Re}(w)^{2}+\operatorname{Im}(w)^{2}+x^{2}}.$
A straightforward calculation then yields
$|h_{-x}(w)|=\frac{|w|}{|w+x|}\leq\frac{|w|}{\sqrt{|w|^{2}+x^{2}}}:=g_{x}(w).$
The function $g_{x}(w)$ is increasing with respect to $|w|$. Noting that $H_{k}=Q_{k}^{*}MQ_{k}$, we apply the interlacing inequalities for singular values in Theorem 2 of [34] to obtain:
$\sigma_{\max}(H_{k})\leq\sigma_{\max}(M)\quad\text{and}\quad\sigma_{\min}(H_{k})\geq\sigma_{\min}(M).$
Furthermore, the modulus of the eigenvalues of a matrix is bounded above by its largest singular value, and the modulus of the smallest eigenvalue is bounded below by its smallest singular value. Letting $\mathcal{S}:=[\sigma_{\min}(M),\sigma_{\max}(M)]$, the problem now turns into analyzing the behavior of $g_{x}$ within $\mathcal{S}$,
(11) $\left\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(f;M,\bm{b})\right\|\leq\left(\frac{1}{\pi}\int_{0}^{\infty}x^{1/2}\cdot\left(\prod_{i=1}^{k}\left\|g_{x}\right\|_{\mathcal{S}}\right)\cdot\left\|h_{-x}(M)\right\|\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|.$
We next give the bound for $\left\|h_{-x}(M)\right\|$. For a fixed real number $x>0$, we consider a new contour $\mathcal{C}$ as in Fig. 1, but with $R=2\sigma_{\max}(M)$ and $\delta=\frac{\epsilon C_{M}}{\sqrt{C_{M}^{2}+4}}$, where $C_{M}=\frac{\lambda_{\min}\left(\frac{M+M^{*}}{2}\right)}{\left|\lambda_{\min}\left(i\left(\frac{M-M^{*}}{2}\right)\right)\right|}$. Using the Cauchy integral formula, it follows that
$h_{-x}(M)=-\frac{1}{2\pi i}\int_{\mathcal{C}}h_{-x}(s)(M-sI)^{-1}\,\mathrm{d}s.$
Then, we have:
$\left\|h_{-x}(M)\right\|$
$\leq$ $\frac{1}{2\pi}\int_{\mathcal{C}_{R}}|h_{-x}(s)|\cdot\left\|(M-sI)^{-1}\right\|\,|\mathrm{d}s|-\frac{1}{2\pi}\int_{\mathcal{C}_{\epsilon}}|h_{-x}(s)|\cdot\left\|(M-sI)^{-1}\right\|\,|\mathrm{d}s|$
$+\frac{1}{2\pi}\int_{\mathcal{C}_{\delta}^{+}}|h_{-x}(s)|\cdot\left\|(M-sI)^{-1}\right\|\,|\mathrm{d}s|-\frac{1}{2\pi}\int_{\mathcal{C}_{\delta}^{-}}|h_{-x}(s)|\cdot\left\|(M-sI)^{-1}\right\|\,|\mathrm{d}s|$
$\leq$ $\frac{1}{2\pi}\int_{\mathcal{C}_{R}}|g_{x}(s)|\cdot\left\|(M-sI)^{-1}\right\|\,|\mathrm{d}s|-\frac{1}{2\pi}\int_{\mathcal{C}_{\epsilon}}|g_{x}(s)|\cdot\left\|(M-sI)^{-1}\right\|\,|\mathrm{d}s|$
$+\frac{1}{2\pi}\int_{\mathcal{C}_{\delta}^{+}}|g_{x}(s)|\cdot\left\|(M-sI)^{-1}\right\|\,|\mathrm{d}s|-\frac{1}{2\pi}\int_{\mathcal{C}_{\delta}^{-}}|g_{x}(s)|\cdot\left\|(M-sI)^{-1}\right\|\,|\mathrm{d}s|.$
Since the integral on the small circular arc tends to 00 as $\epsilon\to 0$, we only need to take care of the integral along $\mathcal{C}^{+}_{\delta}$, $\mathcal{C}^{-}_{\delta}$ and over the large circular arc $\mathcal{C}_{R}$. On the large circular arc $\mathcal{C}_{R}$, we start with the expression:
$\left\|(M-sI)^{-1}\right\|$ $=\frac{1}{\sqrt{\lambda_{\min}\left((M-sI)^{*}(M-sI)\right)}}$
$=\frac{1}{\sqrt{\lambda_{\min}\left(M^{*}M-s^{*}M-sM^{*}+|s|^{2}I\right)}}$
(12) $\leq\frac{1}{\sqrt{\lambda_{\min}\left(M^{*}M\right)+\lambda_{\min}\left(|s|^{2}I\right)+\lambda_{\min}\left(-s^{*}M-sM^{*}\right)}}.$
Letting $T=-s^{*}M-sM^{*}$, we get
$\bm{x}^{*}T\bm{x}$ $=-s^{*}\bm{x}^{*}M\bm{x}-s\bm{x}^{*}M^{*}\bm{x}$
$=-s^{*}\bm{x}^{*}M\bm{x}-s(\bm{x}^{*}M\bm{x})^{*}.$
Thus, one obtains $\lambda_{\min}(T)=\min\limits_{\|\bm{x}\|=1}\bm{x}^{*}T\bm{x}=-2\max\limits_{\|\bm{x}\|=1}\operatorname{Re}(s^{*}\bm{x}^{*}M\bm{x}).$ In addition, applying the Cauchy-Schwarz inquality leads to
$\operatorname{Re}(s^{*}\bm{x}^{*}M\bm{x})\leq|s^{*}\bm{x}^{*}M\bm{x}|=|s|\cdot|\bm{x}^{*}M\bm{x}|\leq|s|\cdot\|M\|=|s|\sigma_{\max}(M).$
Then, we find $\lambda_{\min}(T)\geq-2|s|\sigma_{\max}(M)$. Substituting the inequality into (3.9), we have
$\left\|(M-sI)^{-1}\right\|$ $\leq\frac{1}{\sqrt{(\sigma_{\min}(M))^{2}+|s|^{2}-2|s|\sigma_{\max}(M)}}$
$=\frac{1}{\sigma_{\min}(M)}.$
Thus, over the the large circular arc $\mathcal{C}_{R}$, we obtain the following estimate:
$\frac{1}{2\pi}\int_{\mathcal{C}_{R}}|g_{x}(s)|\cdot\left\|(M-sI)^{-1}\right\||\mathrm{d}s|\leq$ $\frac{2\pi R}{2\pi\sigma_{\min}(M)}\frac{R}{\sqrt{R^{2}+x^{2}}}$
$=$ $\frac{4\sigma_{\max}(M)\kappa(M)}{(\sqrt{4(\sigma_{\max}(M))^{2}+x^{2}})}.$
Along the lines $\mathcal{C}^{+}_{\delta}$, $\mathcal{C}^{-}_{\delta}$, one arrives at
$\left\|(M-sI)^{-1}\right\|$ $=\frac{1}{\sqrt{\lambda_{\min}\left(M^{*}M-s^{*}M-sM^{*}+|s|^{2}I\right)}}$
(13) $\leq\frac{1}{\sqrt{\lambda_{\min}\left(M^{*}M\right)+\lambda_{\min}(-s^{*}M-sM^{*})}}.$
Given that $M$ is positive-definite, $\frac{M+M^{*}}{2}$ is also positive-definite. We assume $s$ takes the form: $s=s_{1}\pm i\delta$. Since $\delta=\frac{\epsilon C_{M}}{\sqrt{C_{M}^{2}+4}}$ with $C_{M}=\frac{\lambda_{\min}\left(\frac{M+M^{*}}{2}\right)}{\left|\lambda_{\min}\left(i\left(\frac{M-M^{*}}{2}\right)\right)\right|}$, then $-R\leq s_{1}\leq-\frac{2\epsilon}{\sqrt{C_{M}^{2}+4}}<-\epsilon$. Therefore,
$\lambda_{\min}(-s^{*}M-sM^{*})$ $=-2\lambda_{\max}\left(s_{1}\left(\frac{M+M^{*}}{2}\right)-i\delta\left(\frac{M-M^{*}}{2}\right)\right)$
$\geq-2\lambda_{\max}\left(s_{1}\left(\frac{M+M^{*}}{2}\right)\right)-2\lambda_{\max}\left(-i\delta\left(\frac{M-M^{*}}{2}\right)\right)$
$=-2s_{1}\lambda_{\min}\left(\frac{M+M^{*}}{2}\right)+2\delta\lambda_{\min}\left(i\left(\frac{M-M^{*}}{2}\right)\right)$
$\geq 0.$
Substituting this into 3.9 yields
$\left\|(M-sI)^{-1}\right\|\leq\frac{1}{\sigma_{\min}(M)}.$
Analogously, with $\delta=\frac{\epsilon C_{M}}{\sqrt{C_{M}^{2}+4}}$, we get
$\frac{1}{2\pi}\int_{\mathcal{C}_{\delta}^{+}}|g_{x}(s)|\cdot\left\|(M-sI)^{-1}\right\|\,|\mathrm{d}s|-\frac{1}{2\pi}\int_{\mathcal{C}_{\delta}^{-}}|g_{x}(s)|\cdot\left\|(M-sI)^{-1}\right\|\,|\mathrm{d}s|$
$\leq$ $\frac{1}{2\pi}\lim_{\begin{subarray}{c}\epsilon\to 0\end{subarray}}\int_{-R}^{-\epsilon}|g_{x}(s_{1}+\delta i)|\cdot\left\|(M-(s_{1}+\delta i)I)^{-1}\right\|\mathrm{d}s_{1}$
$-\frac{1}{2\pi}\lim_{\begin{subarray}{c}\epsilon\to 0\end{subarray}}\int^{-R}_{-\epsilon}|g_{x}(s_{1}-\delta i)|\cdot\left\|(M-(s_{1}-\delta i)I)^{-1}\right\|\mathrm{d}s_{1}$
$=$ $\frac{1}{\pi}\int_{0}^{R}|g_{x}(-s_{1})|\cdot\left\|(M+s_{1}I)^{-1}\right\|\mathrm{d}s_{1}$
$\leq$ $\frac{R^{2}}{\pi\sigma_{\min}(M)(\sqrt{R^{2}+x^{2}})}$
$=$ $\frac{4\sigma_{\max}(M)\kappa(M)}{\pi(\sqrt{4(\sigma_{\max}(M))^{2}+x^{2}})}.$
Finally, combining all estimates, we obtain the upper bound:
(14) $\left\|h_{-x}(M)\right\|\leq\frac{1+\pi}{\pi}\frac{4\sigma_{\max}(M)\kappa(M)}{(\sqrt{4(\sigma_{\max}(M))^{2}+x^{2}})}\leq\frac{4(1+\pi)\kappa(M)}{\pi}\frac{\sigma_{\max}(M)}{(\sqrt{(\sigma_{\max}(M))^{2}+x^{2}})}.$
Step 3: Apply Beta and Gamma Functions. Substituting (14) into (11) gives
$\left\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(f;M,\bm{b})\right\|$
$\leq\left(\frac{1}{\pi}\int_{0}^{\infty}x^{1/2}\cdot\left\|g_{x}(M)\right\|_{\mathcal{S}}^{k}\cdot\left\|h_{-x}\right\|\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|$
$\leq\frac{4(1+\pi)\kappa(M)}{\pi^{2}}\left(\int_{0}^{\infty}x^{1/2}\cdot\left\|g_{x}\right\|_{\mathcal{S}}^{k}\cdot\frac{\sigma_{\max}(M)}{(\sqrt{(\sigma_{\max}(M))^{2}+x^{2}})}\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|$
$\leq\frac{4(1+\pi)\kappa(M)}{\pi^{2}}\left(\int_{0}^{\infty}x^{1/2}\cdot\frac{(\sigma_{\max}(M))^{k+1}}{(\sqrt{(\sigma_{\max}(M))^{2}+x^{2}})^{k+1}}\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|.$
Define $\mathrm{I}:=\int_{0}^{\infty}x^{1/2}\cdot\frac{(\sigma_{\max}(M))^{k+1}}{(\sqrt{(\sigma_{\max}(M))^{2}+x^{2}})^{k+1}}\,\mathrm{d}x$. Let $x=\sigma_{\max}(M)y$, then we obtain
$\rm{I}$ $=(\sigma_{\max}(M))^{3/2}\int^{\infty}_{0}\frac{y^{1/2}}{(\sqrt{1+y^{2}})^{k+1}}\mathrm{~{}d}y$
$=(\sigma_{\max}(M))^{3/2}\int^{\frac{\pi}{2}}_{0}\frac{(\tan(\theta))^{1/2}(\sec(\theta))^{2}}{(\sec(\theta))^{k+1}}\mathrm{~{}d}\theta$
$=(\sigma_{\max}(M))^{3/2}\int^{\frac{\pi}{2}}_{0}(\sin(\theta))^{1/2}(\cos(\theta))^{k-3/2}\mathrm{~{}d}\theta$
$=\frac{(\sigma_{\max}(M))^{3/2}}{2}\mathrm{B}\left(\frac{3}{4},\frac{k}{2}-\frac{1}{4}\right)$
$=\frac{(\sigma_{\max}(M))^{3/2}}{2}\frac{\Gamma(3/4)\Gamma(k/2-1/4)}{\Gamma(k/2+1/2)}.$
Here, $\text{B}(\cdot,\cdot)$ and $\Gamma(\cdot)$ are the standard Beta and Gamma functions, respectively, with $\mathrm{B}(x,y)=\frac{\Gamma(x)\Gamma(y)}{\Gamma(x+y)}.$ Furthermore, noting that in [36] for $w>0$ and $l\in(0,1)$,
$\left(\frac{w}{w+l}\right)^{1-l}\leq\frac{\Gamma(w+l)}{w^{l}\Gamma(w)}\leq 1,$
we arrive at
$\frac{\Gamma(k/2-1/4)}{\Gamma(k/2+1/2)}$ $\leq\frac{(k/2+1/2)^{1/4}}{(k/2-1/4)^{1/4}(k/2-1/4)^{3/4}}$
$\leq\frac{2^{5/4}}{(k-1/2)^{3/4}}$
$\leq\frac{\sqrt{2}\pi^{2}}{(1+\pi)\Gamma(3/4)(k-1/2)^{3/4}}.$
Then, it holds that
$\mathrm{I}\leq\frac{\pi^{2}(\sigma_{\max}(M))^{3/2}}{\sqrt{2}(1+\pi)(k-1/2)^{3/4}}.$
Collecting all estimates leads to
$\left\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(f;M,\bm{b})\right\|\leq\frac{2\sqrt{2}\kappa(M)(\sigma_{\max}(M))^{3/2}}{(k-1/2)^{3/4}}\left\|\bm{\xi}_{0}^{k}\right\|.$
Thus, the relative error of the Arnoldi approximation for matrix square roots is bounded by:
$\frac{\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};M,\bm{b})\|}{\|M^{1/2}\bm{b}\|}$ $\leq\frac{2\sqrt{2}\kappa(M)(\sigma_{\max}(M))^{3/2}}{(k-1/2)^{3/4}}\frac{\left\|\bm{\xi}_{0}^{k}\right\|}{\|M^{1/2}\bm{b}\|}$
(15) $\leq\frac{2\sqrt{2}(\kappa(M))^{5/2}}{(k-1/2)^{3/4}}\frac{\|\bm{r}^{k}_{0}\|}{\|\bm{b}\|}.$
Step 4: Bound $\|\bm{r}^{k}_{0}\|$. From (6), it holds that
$\bm{r}^{k}_{0}$ $=\bm{b}-MQ_{k}H_{k}^{-1}\|\bm{b}\|\bm{e}_{1}$
$=\bm{b}-(Q_{k}H_{k}+h_{k+1,k}\bm{q}_{k+1}\bm{e}_{k}^{T})H_{k}^{-1}\|\bm{b}\|\bm{e}_{1}$
$=-h_{k+1,k}\bm{q}_{k+1}\bm{e}_{k}^{T}H_{k}^{-1}\|\bm{b}\|\bm{e}_{1}.$
Taking the 2-norm on both sides gives
(16) $\left\|\bm{r}^{k}_{0}\right\|\leq h_{k+1,k}\left\|H_{k}^{-1}\right\|\left\|\bm{b}\right\|\leq\frac{h_{k+1,k}}{\sigma_{\min}\left(H_{k}\right)}\left\|\bm{b}\right\|.$
Multiplying $\bm{e}_{k}$ on both sides of (6) and then taking the 2-norm, we have
(17) $h_{k+1,k}=\|M\bm{q}_{k}-Q_{k}H_{k}\bm{e}_{k}\|\leq\|M\bm{q}_{k}\|+\|H_{k}\bm{e}_{k}\|\leq\sigma_{\max}(M)+\sigma_{\max}(H_{k}),$
where the second inequality is achieved by utilizing the triangle inequality and the orthogonality of $Q_{k}$. Plugging the above estimate into (16) and utilizing the interlacing inequalities for singular values, we further bound $\left\|\bm{r}_{0}^{k}\right\|$ by
$\left\|\bm{r}^{k}_{0}\right\|\leq\frac{\sigma_{\max}(M)+\sigma_{\max}(H_{k})}{\sigma_{\min}\left(H_{k}\right)}\left\|\bm{b}\right\|\leq\frac{2\sigma_{\max}\left(M\right)}{\sigma_{\min}\left(M\right)}\left\|\bm{b}\right\|=2\kappa(M)\|\bm{b}\|.$
Consequently, substituting the above estimate of $\bm{r}_{0}^{k}$ into (3.9), we obtain
$\frac{\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};M,\bm{b})\|}{\|M^{1/2}\bm{b}\|}\leq\frac{2\sqrt{2}(\kappa(M))^{5/2}}{(k-1/2)^{3/4}}\frac{\|\bm{r}^{k}_{0}\|}{\|\bm{b}\|}\leq\frac{4\sqrt{2}(\kappa(M))^{7/2}}{(k-1/2)^{3/4}}.$
Remark 3.10.
If the matrix $M$ is Hermitian, then using the results from [6] and [1], we obtain
$\frac{\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};M,\bm{b})\|}{\|M^{1/2}\bm{b}\|}\leq\frac{(\kappa(M))^{3/2}}{2k^{3/2}}\frac{\|\bm{r}^{k}_{0}\|}{\|\bm{b}\|}\leq\frac{(\kappa(M))^{5/2}}{k^{3/2}}.$
This sharper result stems from the fact that the Hermitian matrix is unitarily diagonalizable with real eigenvalues, which allows for a more refined spectral analysis of $\left|\operatorname{det}\left(h_{z}\left(H_{k}\right)\right)\right|$ and $\left\|h_{z}(M)\right\|$ in 3.9.
Remark 3.11.
Note that the residual $\bm{r}_{0}^{k}$ is precisely the residual from the FOM applied to the linear system $M\bm{x}=\bm{b}$. Since the error in Theorem 3.8 is bounded in terms of $\bm{r}_{0}^{k}$, which is naturally computed in the Arnoldi iteration, we thus have a practical stopping criterion for our algorithm.
Remark 3.12.
The bound in (17) is a loose upper bound for $h_{k+1,k}$. In practice, $h_{k+1,k}$ tends to decrease with $k$, as the Krylov subspace increasingly approximates the spectral subspace associated with the dominant eigenvalues of $M$. However, this behavior is not guaranteed in general. To further understand the behavior of $\|\bm{r}_{0}^{k}\|$, one may explore results from [7]. They show that $\min_{0\leq j\leq k}\|\mathbf{r}_{0}^{j}\|\leq\sqrt{k+1}\cdot\|\mathbf{r}^{k}_{\mathrm{G}}\|$, where $\mathbf{r}^{k}_{\mathrm{G}}$ denotes the GMRES residual after $k$ iterations for solving $M\bm{x}=\bm{b}$. This implies that the overall convergence of FOM is at most a factor of $\sqrt{k+1}$ worse than that of GMRES. However, the residual norm at a fixed iteration $k$ can be arbitrarily large, and thus this result does not apply to our setting.
3.3 Preprocessing
In practice, matrices arising from large-scale problems are often too large to be directly measured, stored, or analyzed, necessitating efficient preprocessing. To further reduce the cost, one direction is to employ data-sparse approximations, such as hierarchical or low-rank representations [5, 8, 10, 17, 22, 23]. These methods accelerate matrix-vector multiplications by compressing the underlying structure of the matrix. Under this framework, we modify the results from Theorem 3.8 to get an error bound for approximating the square root of a preprocessed matrix using the Arnoldi iteration. This is summarized in the following theorem.
Theorem 3.13.
Given a positive-definite matrix $M\in\mathbb{C}^{n\times n}$ and a vector $\bm{b}\in\mathbb{C}^{n}$, assume that M admits a data-sparse approximation $\widetilde{M}$ which is positive definite and satisfies, for some $0<\epsilon<1/\kappa(M)$, $\|M-\widetilde{M}\|_{*}\leq\epsilon\|M\|$. Then, the relative error of the Arnoldi process for approximating the matrix square root is bounded by
$\frac{\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};\widetilde{M},\bm{b})\|}{\|M^{1/2}\bm{b}\|}\leq\epsilon^{1/2}\kappa^{1/2}(M)+\frac{4\sqrt{2}(1+\epsilon)^{7/2}(\kappa(M))^{9/2}}{(k-1/2)^{3/4}(1-\epsilon\kappa(M))^{9/2}}.$
Proof 3.14.
Based on $\widetilde{M}$, we apply the Arnoldi process to compute the matrix square root. Now the problem turns into:
$\frac{\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};\widetilde{M},\bm{b})\|}{\|M^{1/2}\bm{b}\|}\leq\frac{\|M^{1/2}\bm{b}-\widetilde{M}^{1/2}\bm{b}\|}{\|M^{1/2}\bm{b}\|}+\frac{\|\widetilde{M}^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};\widetilde{M},\bm{b})\|}{\|M^{1/2}\bm{b}\|}.$
For the first term on the right hand side, according to the Powers–Størmer inequality [32], we have
$\frac{\|M^{1/2}\bm{b}-\widetilde{M}^{1/2}\bm{b}\|}{\|M^{1/2}\bm{b}\|}\leq\frac{\|M^{1/2}-\widetilde{M}^{1/2}\|\|\bm{b}\|}{\|M^{1/2}\bm{b}\|}\leq\frac{\|M-\widetilde{M}\|_{*}^{1/2}\|\bm{b}\|}{\|M^{1/2}\bm{b}\|}\leq\epsilon^{1/2}\kappa^{1/2}(M).$
With Theorem 3.8, the second term is bounded by:
$\frac{\|\widetilde{M}^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};\widetilde{M},\bm{b})\|}{\|\widetilde{M}^{1/2}\bm{b}\|}\leq\frac{4\sqrt{2}(\kappa(\widetilde{M}))^{7/2}}{(k-1/2)^{3/4}}.$
By Weyl’s inequality [4], we have $|\sigma_{k}(M)-\sigma_{k}(\widetilde{M})|\leq\|M-\widetilde{M}\|\leq\|M-\widetilde{M}\|_{*}\leq\epsilon\|M\|=\epsilon\sigma_{\max}(M)$. Rearranging terms gives
$\kappa(\widetilde{M})\leq\frac{(1+\epsilon)\sigma_{\max}(M)}{\sigma_{\min}(M)-\epsilon\sigma_{\max}(M)}=\frac{(1+\epsilon)\kappa(M)}{1-\epsilon\kappa(M)}.$
Therefore, we conclude that
$\frac{\|\widetilde{M}^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};\widetilde{M},\bm{b})\|}{\|M^{1/2}\bm{b}\|}$ $\leq\frac{4\sqrt{2}(\kappa(\widetilde{M}))^{7/2}}{(k-1/2)^{3/4}}\frac{\|M^{1/2}\bm{b}\|}{\|\widetilde{M}^{1/2}\bm{b}\|}$
$\leq\frac{4\sqrt{2}(1+\epsilon)^{7/2}(\kappa(\widetilde{M}))^{9/2}}{(k-1/2)^{3/4}(1-\epsilon\kappa(M))^{9/2}}.$
4 Numerical experiments and discussion
In this section, we report on numerical studies of the Arnoldi iteration for computing the matrix square root applied to a vector, that is, $M^{1/2}\bm{b}$. First, we explore various matrix settings to study the dependence of relative error on the condition number of matrices and the iteration count. We start with a Hermitian matrix, confirming the result in Remark 3.10 as well as Lemma 3.3 in [1]. Then, we test both sparse and general non-Hermitian matrices validating Theorem 3.8. Finally, we explore the use of the Arnoldi iteration for computing $M^{1/2}\bm{b}$ in the simulation of particulate suspensions to demonstrate the effectiveness of the proposed method in practical applications. Consistently across examples, the Arnoldi approximation is terminated based on the relative residual norm for solving $M\bm{x}=\bm{b}$, with a tolerance of $10^{-2}$. In particular, all residuals and relative errors reported in this section are measured in the 2-norm.
4.1 2D Laplacian
In this example, we consider the SPD matrix $M$ arising from a five-point stencil discretization of the Laplace equation with homogeneous Dirichlet boundary conditions on the unit square domain. The discrete problem is formulated on a uniform numerical grid with grid spacing $h=\frac{1}{n}$, where $n$ denotes the number of divisions along each axis. Then, $M$ takes the form of a block-structured matrix:
$M:=\begin{pmatrix}C&D&0&\cdots&0\\ D&C&D&\cdots&0\\ 0&D&C&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&D&C\end{pmatrix}$
where, $C$ and $D$ are defined as:
$C=\operatorname{tridiag}\left(-\frac{1}{h^{2}},\;\frac{4}{h^{2}},\;-\frac{1}{h^{2}}\right)_{(n-1)\times(n-1)},$
$D=\operatorname{diag}\left(-\frac{1}{h^{2}},-\frac{1}{h^{2}},\dots,-\frac{1}{h^{2}}\right)_{(n-1)\times(n-1)}.$
In this case, vector $\bm{b}$ is taken to be the vector of all ones. The performance of the Arnoldi iteration for computing $M^{1/2}\bm{b}$ is presented in Table 1. The numerical experiments confirm that as the condition number of the matrix increases, maintaining a relative error on the order of $10^{-5}$ requires a larger number of iterations. The decay of the normalized residual for FOM and the relative error for the Arnoldi method are depicted in Fig. 2. In this example, a normalized residual of $10^{-2}$ for FOM yields a relative error of approximately $10^{-5}$ in the Arnoldi approximation, showing a three order of magnitude improvement. This observation indicates that employing the normalized residual of FOM as a stopping criterion is both effective and efficient, enabling accurate results even with a relatively large tolerance. To gain further insight into this example, we study the relationship between the relative error, condition number, and iteration count. The log-log plot of the relative error decay against the iteration count $k$ for different condition numbers is shown in Fig. 3. We observe that the error exhibits a decay rate of approximately $\mathcal{O}(k^{-3/2})$ with respect to the iteration count $k$. Additionally, we fix $k=25$ and plot the relative error against the condition number on a log-log scale in Fig. 4. As shown, the error decay scales with the condition number $\kappa$ at a rate close to $\mathcal{O}(\kappa^{5/2})$. These results are consistent with the conclusions for Hermitian matrices stated in Remark 3.10 and Lemma 3.3 in [1].
Table 1: Condition number of $M$, iteration counts, and relative error of the Arnoldi iteration for the discrete 2D Laplacian.
$n$ $\kappa(M)$ Relative Error Iteration Counts $k$
30 529.97 1.90e-05 29
40 942.53 1.59e-05 39
50 1472.96 1.07e-05 50
60 2121.27 9.98e-06 60
70 2887.45 7.84e-06 71
80 3771.51 7.57e-06 81
90 4773.44 6.31e-06 92
100 5893.24 6.22e-06 102
110 7130.92 4.69e-06 114
Figure 2: Semi-log plot of the normalized residual for FOM (*–) and the relative error for Arnoldi (o–) with different condition number against iteration count $k$ for the discrete 2D Laplacian.
Figure 3: Log-log plot of relative error decay with varying condition numbers against iteration count $k$ for the discrete 2D Laplacian.
Figure 4: Log-log plot of relative error with $k=25$ against condition number for the discrete 2D Laplacian.
4.2 Convection-Diffusion problem
In this case, $M$ is taken from the discretization of the convection-diffusion problem, which describes transport phenomena in a given domain. The continuous convection-diffusion operator is defined as
$Lu:=\eta u^{{\prime}{\prime}}(x)+u^{\prime}(x),\quad x\in[0,1],$
subject to homogeneous Dirichlet boundary conditions. Here, $\eta$ is the diffusion coefficient. For small values of $\eta$, the problem is convection-dominated. To accurately capture the convection effects while maintaining numerical stability, we employ an upwind scheme. A uniform numerical grid with grid spacing $h=\frac{1}{n}$ is adopted, where the discrete points are defined as $x_{i}=ih$, $i=0,1,2,\dots,n$. The resulting finite-difference scheme of the operator takes the following form:
$\eta\frac{u_{i+1}-2u_{i}+u_{i-1}}{h^{2}}+\frac{u_{i}-u_{i-1}}{h},$
where $u_{i}$ represents the numerical approximation to $u(x_{i})$. Let $M$ be the tridiagonal matrix obtained from the above finite-difference discretization of $L$. We take $\bm{b}$ to be the vector with all entries equal to one, and fix $\eta$ = 0.1.
The results for approximating $M^{1/2}\bm{b}$ are summarized in Table 2 and Fig. 5, and the general trends align with the conclusions drawn form the 2D Laplacian example. However, the matrix in this case is no longer symmetric. From the plot of the error versus iteration count in Fig. 6, we see that as the condition number increases, the convergence rate closely follows a $k^{-3/4}$ relationship with respect to the iteration count. This observation is consistent with the theoretical result established in Theorem 3.8. Meanwhile, for a fixed $k=480$, Fig. 6 indicates that for large condition numbers, the observed trend closely follows $\kappa^{7/2}$.
Table 2: Condition number of $M$, iteration counts, and relative error of the Arnoldi iteration for the convection-diffusion problem.
$n$ $\kappa(M)$ Relative Error Iteration Counts $k$
500 67510.90 3.99e-08 495
600 97087.72 3.31e-08 594
700 132022.80 2.85e-08 693
800 172316.14 2.53e-08 792
900 217967.73 2.29e-08 891
1000 268977.57 2.10e-08 990
1100 325345.68 1.94e-08 1089
1200 387072.03 1.81e-08 1188
1300 454156.65 1.71e-08 1287
Figure 5: Semi-log plot of the normalized residual for FOM (*–) and the relative error for Arnoldi (o–) with different condition number against iteration count $k$ for the convection-diffusion problem.
Figure 6: Log-log plot of relative error decay with varying condition numbers against iteration count $k$ for the convection-diffusion problem.
Figure 7: Log-log plot of relative error with $k=480$ against condition number for the convection-diffusion problem.
4.3 General matrices
Next, we consider positive definite but non-symmetric dense matrices. The matrix $M$ is constructed as follows:
(18) $M=M_{0}+\alpha I=USV^{T}+\alpha I,$
where $M_{0}$ is a low-rank matrix with rank $m$, $U$ and $V$ are orthonormal matrices generated randomly, and $S$ is an $m\times m$ diagonal matrix whose entries decay logarithmically from 1 to $\frac{1}{\beta}$, where $\beta$ is a given constant. A shift term, $\alpha I$, is introduced to ensure positive-definiteness of $M$. Specifically, the shift parameter $\alpha$ is defined as: $\alpha=-(1+5\times 10^{-3})\min\operatorname{Re}(\lambda_{\min}(M_{0})).$ As a result, the condition number of the $M$ is influenced by both $\alpha$ and $\beta$. In this test, we set the rank to $m=500$, and the size of the matrix $M$ is $5000\times 5000$. The parameter $\beta$ takes values of 400, 600, 800, 1000, 1200, 1400, 1600, 1800, 2000. As before, $\bm{b}$ is taken to be the vector of all ones.
In Table 3, we display the corresponding results. As the relative error remains around $10^{-5}$, the number of iterations generally increases with the condition number. Similarly, the normalized residual for solving $M\bm{x}=\bm{b}$ provides an effective stopping criterion for the Arnoldi iteration in the evaluation of $M^{1/2}\bm{b}$ (Fig. 8). As seen in Fig. 9, the error decay of the Arnoldi method closely follows the reference slope before the transition to superlinear convergence. However, we observe that with the iteration count fixed at $k=35$, the fitted slope is approximately $0.9$, which is much smaller than the theoretical value of $7/2$. These observations indicate that the theoretical results established in Theorem 3.13 yield a reasonably tight estimate for the iteration count $k$, but only a loose upper bound with respect to the condition number for certain type of matrices. This may be due to the fact that the bounds could depend not only on the condition number and iteration counts but also on intrinsic spectral characteristics or a specific structure of the matrices.
Table 3: Condition number of $M$, iteration counts, and relative error of the Arnoldi iteration for general matrices constructed via Eq. 18.
$\kappa(M)$ Relative Error Iteration Counts $k$
2192.38 2.40e-05 38
5375.35 2.90e-05 41
9325.42 2.62e-05 56
28826.11 3.86e-05 38
32437.66 2.45e-05 47
47395.74 2.87e-05 58
52845.93 6.31e-05 36
166131.16 5.27e-05 63
188372.52 1.03e-04 51
Figure 8: Semi-log plot of the normalized residual for FOM (*–) and the relative error for Arnoldi (o–) with different condition number against iteration count $k$ for general matrices constructed via Eq. 18.
Figure 9: Log-log plot of relative error decay with varying condition numbers against iteration count $k$ for general matrices constructed via Eq. 18.
Figure 10: Log-log plot of relative error with $k=35$ against condition number for general matrices constructed via Eq. 18.
4.4 Practical application
The study of particulate suspensions is of great interest, and plays critical roles in diverse fields such as materials design [27], biological systems [27], micro-robotics [37], and food processing [21]. Immersed in a fluid environment, particles’ mobility is dictated by the hydrodynamic interactions between the particles. At micro scales, particles are also subject to thermal fluctuations, in the form of Brownian motion. The configuration of particles evolves over time and is governed by the overdamped limit of Langevin dynamics [14]. Thus, the computation of particles’ Brownian displacements crucially rely on the mobility matrix, and its square root, both dependent on all particles’ instantaneous positions. Efficiently computing these two matrices remains a significant computational challenge.
Since hydrodynamic interactions depend on the separation distance between particles, the particle interactions weaken as the distance increases [25]. As a result, far-field blocks in the mobility matrix exhibit low-rank structure, and can then be efficiently approximated. To this end, the Hydrodynamic Interaction Graph Neural Network (HIGNN) framework, proposed in [25, 24], maps the many-body hydrodynamic interactions onto a graph neural network. By employing hierarchical-matrices using adaptive cross approximation [17], the computational complexity for using HIGNN to compute particles’ velocities given any applied forces can be reduced to nearly linear [26].
In the following tests, $M$ is the mobility matrix obtained from the HIGNN model and has been preprocessed into a hierarchical-matrix form. Also, we assume that a uniform gravitational force $\bm{b}$ is applied to each particle. This numerical example focuses on evaluating the effectiveness of the Arnoldi iteration on computing $M^{1/2}$ in a long-time dynamic simulation of particles’ sedimentation. In the simulations, the explicit Euler method was adopted as the temporal integrator to update particles’ positions from their velocities, with a time step of $\Delta t=0.005$. At each time step, the HIGNN is applied to predict the velocities of all particles. To generate the initial cubic-lattice like configuration of particles, the particles are first positioned at grid points with a spacing of 4, and then each particle is randomly perturbed with a displacement of magnitude 0.2 from its original position. With this initial configuration, we simulate particles sedimenting in an unbounded domain. The number of particles, $N$, considered here varies from $20^{3}$ to $100^{3}$. For such large systems, direct computation of the square root of the mobility matrix $M$ can be computationally expensive. To address this and evaluate the accuracy of the Arnoldi iteration in this context, we apply the Arnoldi process twice to approximate $M\bm{b}=M^{1/2}(M^{1/2}\bm{b})$. Thus, the relative error is defined as $\frac{\|M\bm{b}-\operatorname{Arn}_{k}(x^{1/2};M,\operatorname{Arn}_{k}(x^{1/2};M,\bm{b}))\|}{\|M\bm{b}\|}$.
Figure 11: Snapshots of 8000 particles sedimenting in an unbounded domain at times $t=0.005$, $0.25$, and $1$, respectively, simulated using the HIGNN. The color is correlated to the particles’ velocity magnitude.
Table 4: Arnoldi Approximation for lattice structure system with constant particle concentration.
$N$ Iteration Counts $k$ Relative Error
$20^{3}$ 10 2.06699e-04
$40^{3}$ 14 1.34295e-04
$60^{3}$ 15 1.70283e-04
$80^{3}$ 18 3.11680e-04
$100^{3}$ 21 3.86345e-04
The snapshots at the 1st, 100th, and 200th time steps from the simulation of a $20^{3}$ particle system are presented in Fig. 11. Letting $M$ be the mobility matrix at the 100th time step, Table 4 records the number of iterations $k$ required to compute $M^{1/2}\bm{b}$ and the relative error as the number of particles $N$ increases from $20^{3}$ to $100^{3}$. Note that in this scenario, since the initial spacing between particles is kept the same as $N$ increases, a consistent particle concentration is maintained. One sees that as the number of particles increases, the number of iterations required to compute the matrix square root applied to a vector remains relatively low and does not grow significantly, while maintaining a relative error on the order of $10^{-4}$. This is because the matrix is well conditioned. Furthermore, we depict the relationship between the number of particles and the runtime on a log-log scale in Fig. 12. The plot implies that the Arnoldi approximation, based on the hierarchical matrix representation, exhibits an $O(N\log N)$ time complexity.
Figure 12: Log-log plot of runtime against number of particles for lattice structure system.
Next, we consider $40^{3}$ particles arranged in a similar initial cubic lattice as above, but with the grid spacing $d$ varying from 3 to 5 (still perturbed by a displacement of 0.2). With this change, we consider different particle concentrations; when $d$ increases, the particle concentration decreases, thus lowering the required rank of the approximation to the corresponding blocks. As shown in Table 5, lower particle concentration leads to reduced runtime for the Arnoldi approximation. However, the required number of iterations remains relatively small and stable compared to the matrix size. In this test, $M$ corresponds to the mobility matrix attained at the 50th time step.
Table 5: Arnoldi Approximation for lattice structure with different particle concentration.
$d$ Iteration Counts $k$ relative error runtime
$3$ 12 1.93793e-04 362.4649
$3.5$ 12 1.81240e-04 386.0504
$4$ 10 1.88017e-04 319.9872
$4.5$ 10 1.86118e-04 307.9138
$5$ 10 1.83476e-04 298.9904
5 Concluding remarks
In this work, we derive a priori error estimates of the Arnoldi process for approximating the matrix square root applied to a vector, extending the results in [6] for Hermitian matrices to non-Hermitian cases. Moreover, we provide an error bound based on a data-sparse representation of the matrix. Through numerical experiments, we investigate the convergence behavior of the Arnoldi iteration for computing $M^{1/2}\bm{b}$ across different matrix types. The observations indicate that the theoretical results presented in Theorem 3.8 appear to be sharp in terms of iteration count, yet remain a loose upper bound with respect to the condition number for certain classes of matrices. Additionally, we consider the mobility matrix arising from particulate suspension simulations, where the matrix is approximated into a hierarchical matrix. The numerical results demonstrate the effectiveness of our approach for large-scale problems. Future extensions of this paper include deriving a sharper dependence on the condition number for approximating the matrix square root applied to a vector via Arnoldi iteration, as well as investigating the error behavior of GMRES in the approximation of $f(M)\bm{b}$ for general functions $f$.
Acknowledgments
The authors would like to thank Zhan Ma and Zisheng Ye for their valuable discussions and assistance with the code implementation. The authors also gratefully acknowledge support from the National Science Foundation under Grant No. DMS-2208267 and the Army Research Office under Grant No. W911NF2310256.

### v3

@@
 The Arnoldi process provides an efficient framework for approximating functions of a matrix applied to a vector, i.e., of the form $f(M)\bm{b}$, by repeated matrix-vector multiplications.
-In this paper, we derive an a priori error estimate for approximating the action of a matrix square root using the Arnoldi process, where the integral representation of the error is reformulated in terms of the error for solving the linear system $M\bm{x}=\bm{b}$.
+In this paper, we derive error estimates for approximating the action of a matrix square root using the Arnoldi process, where the integral representation of the error is reformulated in terms of the error for solving the linear system $M\bm{x}=\bm{b}$.
 The results extend the error analysis of the Lanczos method for Hermitian matrices in [Chen et al., SIAM J.
 Matrix Anal.
-Appl., 2022] to non-Hermitian cases.
-Furthermore, to make the method applicable to large-scale problems, we assume that the matrices are preprocessed utilizing data-sparse approximations preserving positive definiteness, and then establish a refined error bound in this setting.
+Appl., 2022] to non-Hermitian cases and provide an improved bound for the Hermitian case.
+Furthermore, in practical settings, the matrix may only be available via approximate or structured representations.
+Motivated by this, we extend the analysis and establish a generalized error bound for perturbed matrices.
 The numerical results on matrices with different structures demonstrate that our theoretical analysis yields a reliable upper bound.
-Finally, simulations on large-scale matrices arising in particulate suspensions validate the effectiveness and practicality of the approach.
+Finally, simulations on large-scale matrices arising in particulate suspensions, represented in hierarchical matrix form, validate the effectiveness and practicality of the approach.
 keywords:
@@
 1 Introduction
-Given a matrix $M\in\mathbb{C}^{n\times n}$, its square roots are the solutions of $X^{2}=M.$ If $M$ is positive definite, then $M$ admits a unique positive definite square root [19], denoted as $M^{1/2}$.
-Matrix square roots frequently arise in stochastic simulation [12], state estimation [2], image reconstruction [3], and Bayesian optimization [31, 18].
+Given a matrix $M\in\mathbb{C}^{n\times n}$, its square roots are the solutions of $X^{2}=M.$ If $M$ is positive definite, then $M$ admits a unique positive definite square root [higham2008functions], denoted as $M^{1/2}$.
+Matrix square roots frequently arise in stochastic simulation [ColloidsReview_COCIS2019], state estimation [asl2019adaptive], image reconstruction [bardsley2012mcmc], and Bayesian optimization [pleiss2020fast, hernandez2014predictive].
 Hence, the computation of a matrix square root is a fundamental problem in numerical linear algebra.
 There are a list of approaches in the literature for computing the square root of a matrix.
-Cholesky decomposition and eigenvalue decomposition provide useful tools for symmetric positive definite (SPD) matrices [35].
-For general matrices, the Schur decomposition transforms the matrix into a triangular form, where the square root can be computed explicitly [19].
-Additionally, Newton-like iterations [19] provide stable and efficient implementation for approximating the matrix square root.
-Unfortunately, these classical methods all suffer from at least $\mathcal{O}(n^{3})$ complexity for large-scale matrices [19].
-On the other hand, even when $M$ is sparse or structured, its square root is typically dense, making direct computation and storage costly or even infeasible for large-scale problems.
-Consequently, rather than computing the square root of a matrix explicitly, the main interest lies in computing the action of a matrix square root on a vector.
+Cholesky decomposition and eigenvalue decomposition provide useful tools for symmetric positive definite (SPD) matrices [watkins2004fundamentals].
+For general matrices, the Schur decomposition transforms the matrix into a triangular form, where the square root can be computed explicitly [higham2008functions].
+Additionally, Newton-like iterations [higham2008functions] provide stable and efficient implementation for approximating the matrix square root.
+However, these classical methods all suffer from at least $\mathcal{O}(n^{3})$ complexity for large-scale matrices [higham2008functions].
+Even when $M$ is sparse or structured, its square root is typically dense, making direct computation and storage costly or even infeasible for large-scale problems.
+Consequently, rather than computing the square root of a matrix explicitly, one can instead compute the action of a matrix square root on a vector.
 This is also the case for general matrix functions.
@@
 For general dense matrices, the overall computational cost of the Arnoldi process is about $\mathcal{O}(kn^{2})$, where $k$ is the number of Arnoldi iterations.
-When the matrix is Hermitian, the Arnoldi process simplifies to the Lanczos algorithm, characterized by a three-term recurrence.
-Strategies such as restarting [13, 15, 20], truncated orthogonalization [30, 9], and randomized sketching [30, 9, 29, 16] have been explored to reduce memory and orthogonalization cost.
-Recently, there has been growing interest in the theoretical analysis of Krylov methods for approximating the action of matrix functions.
-As shown in [6], the error of the Lanczos algorithm after $k$ iterations for approximating $f(M)\mathbf{b}$ is bounded by the error of the best $k$-degree polynomial that uniformly approximates $f(x)$ over the spectral interval $[\lambda_{\min}(M),\lambda_{\max}(M)]$, where $\lambda_{\min}(M)$ and $\lambda_{\max}(M)$ denote the smallest and largest eigenvalues of the matrix $M$, respectively.
-Such a classical bound has also been studied within the setting of finite precision arithmetic [28].
-In [11], a posteriori error estimates for polynomial Krylov methods are developed and applied to a broad class of matrix functions with integral representations.
-Based on the Cauchy integral formula, Chen et al.
-[6] establish a priori and a posteriori error bounds for the Lanczos method by reformulating the approximation error in terms of the error of solving a shifted linear system.
-In addition, the authors in [1] explore the near optimality of the Lanczos methods for rational functions with real poles lying outside the spectral interval of $M$.
-Their results are established under the assumption that the matrices are Hermitian, while in many practical applications, the matrices of interest are non-Hermitian.
-This motivates us to extend their results to the non-Hermitian setting.
-In this work, we present an a priori error bound of the Arnoldi process specifically for evaluating a matrix square root, extending the results in [6] to the case of positive-definite non-Hermitian matrices.
+When the matrix is Hermitian, the Arnoldi process simplifies to the Lanczos algorithm [chen2024lanczos], characterized by a three-term recurrence.
+There has been growing interest in the study and discussion of Krylov methods for approximating the action of matrix functions.
+An early study by Saad [saad1992analysis] related the error of Krylov subspace approximations for $f(M)\mathbf{b}$ to polynomial approximation.
+If $M$ is Hermitian, then one obtains the near-optimality property of Lanczos, that is, the error after $k$ iterations is bounded by that of the best polynomial of degree less than $k$ uniformly approximating $f(x)$ over the spectrum of $M$; see, e.g.
+[musco2018stability, chen2022error].
+In [diele2009error], the error of polynomial Krylov approximations for matrix functions is represented in integral forms, and a posteriori error estimates for exponential-like and trigonometric matrix functions are provided.
+In [eiermann2006restarted, ilic2010restarted], the authors expressed the errors of Krylov subspace approximations in terms of divided differences, providing insights into the design of restart schemes.
+For general analytic functions $f$ admitting an integral representation, [frommer2014efficient] provided an integral representation for the error of the Arnoldi approximation to $f(M)\bm{b}$, and proposed a restarting strategy based on evaluating this error function via numerical quadrature.
+Based on [frommer2014efficient, frommer2014convergence], Frommer and Schweitzer [frommer2016error] derived error bounds and estimates of the Krylov subspace approximations $f(M)\bm{b}$ for Stieltjes functions.
+Using the Cauchy integral formula, Chen et al.
+[chen2022error] established a priori and a posteriori error bounds for the Lanczos method by reformulating the approximation error in terms of the error of solving a shifted linear system.
+In addition, Amsel et al.
+[amsel2023near] explored the near optimality of the Lanczos methods for rational functions with real poles lying outside the spectral interval of $M$.
+Recently, [schweitzer2025near] proved that, for Stieltjes and related functions, the Lanczos method exhibits, what they call, near-instance optimality.
+There are also several works that studied Krylov approximations in the context of rational approximations [frommer2009error, frommer2008stopping, frommer20132] and matrix exponentials [beckermann2009error, stewart1996error, hochbruck1997krylov, druskin1998using].
+Most of these results have been established under the assumption that the matrices are Hermitian, while in many practical applications, the matrices of interest are non-Hermitian, which motivates us to extend the results to the non-Hermitian setting.
+In this work, we present error bounds of the Arnoldi process specifically for evaluating a matrix square root, extending the results in [chen2022error] to the case of positive-definite non-Hermitian matrices.
+In this context, we say that $M$ is positive-definite if $\operatorname{Re}(\bm{x}^{*}M\bm{x})>0$ for any unit vector $\bm{x}\in\mathbb{C}^{n}$.
 The main idea is to utilize the integral representation of the error in the iterations of the Arnoldi process, and then relate it to the error for solving the linear system $M\bm{x}=\bm{b}$.
 Compared to Hermitian matrices, non-Hermitian matrices exhibit complex spectral behavior: they are not necessarily diagonalizable, and even when they are, the eigenvectors are typically not orthogonal and the eigenvalues may be complex.
-To address these difficulties, we employ two techniques: the first is applying the Cauchy integral formula to estimate the norms of matrix functions for general matrices that may not be diagonalizable; the second is relating the modulus of a function in the complex plane to the real-valued function of the modulus, reducing the analysis to the positive real line.
-Moreover, for large-scale problems, we derive an error bound for approximating the matrix square root based on a data-sparse representation of the matrix.
+To address these difficulties, we relate the modulus of a function in the complex plane to the real-valued function of the modulus, reducing the analysis to the positive real line.
+This also allows us to derive a sharper bound for the Hermitian case by using the average of eigenvalues, which improves on the results in [chen2022error, amsel2023near].
+Additionally, for large-scale problems, we derive an error bound for approximating the matrix square root based on a data-sparse representation of the matrix.
 In numerical experiments, we investigate matrices with different structures to verify the theoretical results.
-Furthermore, we validate the effectiveness of our approach on large-scale mobility matrices arising from particulate suspension simulations.
+Finally, we validate the effectiveness of our approach on large-scale mobility matrices arising from particulate suspension simulations.
 To improve computational efficiency and reduce storage cost, the matrices are preprocessed through adaptive cross approximation and stored in the form of hierarchical matrices.
-The remainder of this paper is arranged as follows.
+The remainder of this paper is organized as follows.
 In Section 2, we introduce the Arnoldi process for matrix functions.
-Error estimation of the Arnoldi process for computing a square root is provided in Section 3.
-Section 4 presents the implementation and reports our numerical results.
-Finally, concluding remarks are made in Section 5.
+The error estimation of the Arnoldi process for computing a square root is provided in Section 3.
+In Section 4, numerical examples are presented to validate the theoretical results discussed in this paper.
+Finally, concluding remarks are provided in Section 5.
 2 Matrix functions and Arnoldi iterations
@@
 $M=ZJZ^{-1},$
-where $Z$ is a nonsingular matrix, and $J=\operatorname{diag}\left(J_{1},J_{2},\ldots,\right.\left.J_{p}\right)$ with $J_{k}\in\mathbb{R}^{m_{k}}$, $\sum_{i=1}^{p}m_{k}=n$ being a Jordan block associated with the eigenvalue $\lambda_{k}$, $1\leq k\leq p$.
+where $Z$ is a nonsingular matrix, and $J=\operatorname{diag}\left(J_{1},J_{2},\ldots,\right.\left.J_{p}\right)$ with $J_{k}\in\mathbb{R}^{m_{k}\times m_{k}}$, $\sum_{i=1}^{p}m_{k}=n$, being a Jordan block associated with the eigenvalue $\lambda_{k}$, $1\leq k\leq p$.
 Then, the matrix function $f(M)$ is defined as
@@
 $f(M)=-\frac{1}{2\pi i}\int_{\mathcal{C}}f(z)(M-zI)^{-1}\,\mathrm{d}z,$
-where $\mathcal{C}$ is a closed contour enclosing the spectrum of $M$. The two definitions are equivalent when $f$ is analytic [19].
+where $\mathcal{C}$ is a closed contour enclosing the spectrum of $M$. The two definitions are equivalent when $f$ is analytic [higham2008functions].
+In practice, we are primarily interested in the action of a matrix function $f(M)$ on a vector $\bm{b}$.
+To achieve this efficiently, we employ the Arnoldi process, which is described in Algorithm 1.
+Given $\bm{b}\in\mathbb{C}^{n}$, this yields
+(6) $MQ_{k}=Q_{k}H_{k}+h_{k+1,k}\bm{q}_{k+1}\bm{e}_{k}^{T},$
+where $H_{k}=[h_{i,j}]\in\mathbb{C}^{k\times k}$ is an upper Hessenberg matrix satisfying $Q_{k}^{*}MQ_{k}=H_{k}$, $\bm{e}_{k}$ is the $k$-th canonical basis vector, and $Q_{k}$ is an orthonormal matrix whose columns, $\bm{q}_{i}$, form a basis for the $k$-th Krylov space,
+$\mathcal{K}_{k}(M,\bm{b})=\text{span}\{\bm{b},M\bm{b},M^{2}\bm{b},\dots,M^{k-1}\bm{b}\}.$
+Then, the $k$-th Arnoldi iteration for approximating $f(M)\bm{b}$ is given by
+$f(M)\bm{b}\approx\operatorname{Arn}_{k}(f;M,\bm{b}):=Q_{k}f(H_{k})Q_{k}^{*}\bm{b}=\|\bm{b}\|Q_{k}f(H_{k})\bm{e}_{1}.$
 Algorithm 1 Arnoldi Process
@@
 15:end for
-In practice, we are primarily interested in the action of a matrix function $f(M)$ on a vector $\bm{b}$.
-To achieve this efficiently, we employ the Arnoldi process, which is described in Algorithm 1.
-Given $\bm{b}\in\mathbb{C}^{n}$, this yields
-(6) $MQ_{k}=Q_{k}H_{k}+h_{k+1,k}\bm{q}_{k+1}\bm{e}_{k}^{T},$
-where $H_{k}=[h_{i,j}]\in\mathbb{C}^{k\times k}$ is an upper Hessenberg matrix satisfying $Q_{k}^{*}MQ_{k}=H_{k}$, $\bm{e}_{k}$ is the $k$-th canonical basis vector, and $Q_{k}$ is an orthonormal matrix whose columns, $\bm{q}_{i}$, form a basis for the $k$-th Krylov space,
-$\mathcal{K}_{k}(M,\bm{b})=\text{span}\{\bm{b},M\bm{b},M^{2}\bm{b},\dots,M^{k-1}\bm{b}\}.$
-Then, the $k$-th Arnoldi iteration for approximating $f(M)\bm{b}$ is given by
-$f(M)\bm{b}\approx\operatorname{Arn}_{k}(f;M,\bm{b}):=Q_{k}f(H_{k})Q_{k}^{*}\bm{b}=\|\bm{b}\|Q_{k}f(H_{k})\bm{e}_{1}.$
-Note that when $f(x)=1/x$, the Arnoldi approximation reduces to the Full Orthogonalization Method (FOM) [33] for solving the linear system $M\bm{x}=\bm{b}$, with initial guess $\bm{x}_{0}=\bm{0}$.
+Note that when $f(x)=1/x$, the Arnoldi approximation reduces to the Full Orthogonalization Method (FOM) [saad2003iterative] for solving the linear system $M\bm{x}=\bm{b}$, with initial guess $\bm{x}_{0}=\bm{0}$.
 For the rest of this paper, whenever FOM is referred, we always assume the initial guess is zero.
 3 Error bound
-In this section, we study an a priori error bound of the Arnoldi process for approximating general matrix functions and then focus on the analysis of the matrix square root.
+In this section, we study error bounds of the Arnoldi process for approximating general matrix functions and then focus on the analysis of the matrix square root.
 3.1 Integral representation of the error
@@
 (7) $f(M)\bm{b}-\operatorname{Arn}_{k}(f;M,\bm{b})=f(M)\bm{b}-\|\bm{b}\|Q_{k}f(H_{k})\bm{e}_{1}=-\frac{1}{2\pi i}\int_{\mathcal{C}}f(z)\bm{\xi}_{z}^{k}\,\mathrm{d}z.$
+Following a similar argument to that in [chen2022error], we have the following results.
 Lemma 3.1.
@@
 Lemma 3.3.
-Assuming $M$ and $M-zI$ are invertible for some $z\in\mathbb{C}$, we obtain
-$\bm{\xi}_{z}^{k}=\det(h_{z}(H_{k}))h_{z}(M)\bm{\xi}_{0}^{k}\quad\text{and}\quad\bm{r}_{z}^{k}=\det(h_{z}(H_{k}))\bm{r}_{0}^{k},$
-where the function $h_{z}(x)$ is defined as $h_{z}(x):=(x-z)^{-1}x$.
+Assuming $M$ and MzIM-zI are invertible for some $z\in\mathbb{C}$, we obtain
+$\bm{\xi}_{z}^{k}=\frac{\det(H_{k})}{\det(H_{k}-zI)}(M-zI)^{-1}M\bm{\xi}_{0}^{k}\quad\text{and}\quad\bm{r}_{z}^{k}=\frac{\det(H_{k})}{\det(H_{k}-zI)}\bm{r}_{0}^{k}.$
 Proof 3.4.
-The proof follows a similar argument to that of Corollary 2.4 in [6].
+The proof follows a similar argument to that of Corollary 2.4 in [chen2022error].
 Then, we have the following Theorem.
@@
 Given a nonsingular matrix $M\in\mathbb{C}^{n\times n}$ and a vector $\bm{b}\in\mathbb{C}^{n}$, let function $f$ be analytic on and inside a closed contour $\mathcal{C}$ enclosing the spectrum of $M$ and $H_{k}$, where $H_{k}$ is an upper Hessenberg matrix generated from the Arnoldi process for $M$ and $\bm{b}$.
-Then, the error of the Arnoldi iteration for $f(M)b$ can be expressed as
-(8) $f(M)\bm{b}-\operatorname{Arn}_{k}(f;M,\bm{b})=\left(-\frac{1}{2\pi i}\int_{\mathcal{C}}f(z)\operatorname{det}\left(h_{z}\left(H_{k}\right)\right)h_{z}(M)\,\mathrm{d}z\right)\bm{\xi}_{0}^{k}.$
+Then, the error of the Arnoldi iteration for $f(M)\bm{b}$ can be expressed as
+(8) $f(M)\bm{b}-\operatorname{Arn}_{k}(f;M,\bm{b})=\left(-\frac{1}{2\pi i}\int_{\mathcal{C}}f(z)\frac{\det(H_{k})}{\det(H_{k}-zI)}(M-zI)^{-1}M\,\mathrm{d}z\right)\bm{\xi}_{0}^{k}.$
 Proof 3.6.
@@
 $f(M)\bm{b}-\operatorname{Arn}_{k}(f;M,\bm{b})$ $=-\frac{1}{2\pi i}\int_{\mathcal{C}}f(z)\bm{\xi}_{z}^{k}\,\mathrm{d}z$
-$=-\frac{1}{2\pi i}\int_{\mathcal{C}}f(z)\det(h_{z}(H_{k}))h_{z}(M)\bm{\xi}_{0}^{k}\,\mathrm{d}z.$
+$=-\frac{1}{2\pi i}\int_{\mathcal{C}}f(z)\frac{\det(H_{k})}{\det(H_{k}-zI)}(M-zI)^{-1}M\bm{\xi}_{0}^{k}\,\mathrm{d}z.$
 Remark 3.7.
-Theorem 3.5 shows that for any nonsingular $M$, the error of the Arnoldi iteration for computing $f(M)\bm{b}$ is reformulated as the error in solving the linear system $M\bm{x}=\bm{b}$.
+Here we assume that the function $f$ is analytic on and inside a closed contour $\mathcal{C}$ enclosing the spectrum of $M$ and $H_{k}$.
+However, the theorem holds under the sufficient condition that $\mathcal{C}$ encloses the field of values of $M$.
+Similarly to the results in [frommer2009error, frommer2014efficient], Theorem 3.5 shows that for any nonsingular $M$, the error of the Arnoldi iteration for computing $f(M)\bm{b}$ is reformulated as the error in solving the linear system $M\bm{x}=\bm{b}$.
 This further simplifies the contour integral.
-For a more general setting, the error admits a bound related to the error of solving shifted linear systems of the form $(M-wI)\bm{x}=\bm{b}$, where $w$ lies outside the spectrum of both $H_{k}$ and $M$; see Corollary 2.5 in [6] for details.
+For a more general setting, the error admits a bound related to the error of solving shifted linear systems of the form $(M-wI)\bm{x}=\bm{b}$, where $w$ lies outside the spectrum of both $H_{k}$ and $M$.
 3.2 Matrix square root
-We now turn to deriving an a priori error estimate of the Arnoldi process for approximating a matrix square root, that is $f(x)=x^{1/2}$.
-To ensure the existence of a unique positive definite square root, we assume that $M$ is positive definite.
-Here, the matrices of interest are non-Hermitian.
-For such matrices, the eigenvalues may lie off the real axis leading to a contour integral in the complex plane for the error of the Arnoldi process, thereby introducing challenges.
-We derive the error bound starting from Theorem 3.5, and organize the proof into the following steps:
-Step 1: Select an appropriate contour in the complex plane, and derive an initial error bound for the 2-norm of the Arnoldi approximation.
-Step 2: Bound the modulus of the complex function $h_{z}(w)$ using a real-valued function of the modulus, and then estimate $\|h_{z}(M)\|$ with a new contour integral.
-Step 3: Apply the properties of the Beta and Gamma functions to obtain an error bound in terms of the residual $\bm{r}_{0}^{k}$ for solving $Mx=b$.
-Step 4: Conclude the proof by deriving a upper bound for $\bm{r}_{0}^{k}$.
-From this, we establish the following theorem.
+We now turn to deriving error estimates of the Arnoldi process for approximating a matrix square root, that is $f(x)=x^{1/2}$.
+To ensure the existence of a unique positive definite square root, we assume that $M$ is positive-definite.
 Theorem 3.8.
-Given a positive-definite matrix $M\in\mathbb{C}^{n\times n}$ and a vector $\bm{b}\in\mathbb{C}^{n}$, the relative error of the Arnoldi process for approximating the matrix square root is bounded by
-$\frac{\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};M,\bm{b})\|}{\|M^{1/2}\bm{b}\|}\leq\frac{2\sqrt{2}(\kappa(M))^{5/2}}{(k-1/2)^{3/4}}\frac{\|\bm{r}^{k}_{0}\|}{\|\bm{b}\|}\leq\frac{4\sqrt{2}(\kappa(M))^{7/2}}{(k-1/2)^{3/4}},$
-where $\kappa(M)$ denotes the condition number of $M$, defined as $\kappa(M)=\frac{\sigma_{\max}(M)}{\sigma_{\min}(M)}$, and $\sigma_{\max}(M)$ and $\sigma_{\min}(M)$ denote the largest and smallest singular values of $M$, respectively
+Given a positive-definite matrix $M\in\mathbb{C}^{n\times n}$ and a vector $\bm{b}\in\mathbb{C}^{n}$, the error of the Arnoldi process for approximating the matrix square root is bounded by
+$\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};M,\bm{b})\|\leq\frac{1}{\pi}\left(\int_{0}^{\infty}x^{1/2}\prod_{i=1}^{k}\left|\frac{\lambda_{i}(H_{k})}{\lambda_{i}(H_{k})+x}\right|\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|,$
+where $\lambda_{1}(H_{k})\geq\lambda_{2}(H_{k})\geq\cdots\geq\lambda_{k}(H_{k})$ are the eigenvalues of $H_{k}$.
 Proof 3.9.
-Step 1: Choose an Appropriate Contour.
-Define $\mathcal{C}$ as a keyhole contour, consisting of a large circular arc $\mathcal{C}_{R}$ of radius $R$ centered at the origin, a small indented circular arc $\mathcal{C}_{\epsilon}$ of radius $\epsilon$ around the origin, and two horizontal line segments $\mathcal{C}^{+}_{\delta}$ and $\mathcal{C}^{-}_{\delta}$, which run at a small vertical distance $0<\delta<\epsilon$ above and below the branch cut, respectively (see Fig.
+Define $\mathcal{C}$ as a keyhole contour, consisting of a large circular arc $\mathcal{C}_{R}$ of radius $R$ centered at the origin, a small circular arc $\mathcal{C}_{\epsilon}$ of radius $\epsilon$ around the origin, and two horizontal line segments $\mathcal{C}^{+}_{\delta}$ and $\mathcal{C}^{-}_{\delta}$, which run at a small vertical distance $0<\delta<\epsilon$ above and below the branch cut, respectively (see Fig.
 1 for details).
@@
 Figure 1: Keyhole contour with outer circle $\mathcal{C}_{R}$, inner circle $\mathcal{C}_{\epsilon}$, and offset $\delta$ above/below the branch cut.
+Using properties of the determinant, we have
+$\left|\frac{\det(H_{k})}{\det(H_{k}-zI)}\right|=\left|\prod_{i=1}^{k}\frac{\lambda_{i}\left(H_{k}\right)}{\lambda_{i}(H_{k}-zI)}\right|=\prod_{i=1}^{k}\left|\frac{\lambda_{i}\left(H_{k}\right)}{\lambda_{i}(H_{k})-z}\right|.$
 Taking the 2-norm on both sides of (8) with $f(z)=z^{1/2}$, we obtain
-$\left\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(f;M,\bm{b})\right\|\leq$ $\left(\frac{1}{2\pi}\int_{\mathcal{C}_{R}}|z^{1/2}|\cdot\left|\operatorname{det}\left(h_{z}\left(H_{k}\right)\right)\right|\cdot\left\|h_{z}(M)\right\|\,|\mathrm{d}z|\right)\left\|\bm{\xi}_{0}^{k}\right\|$
-$-\left(\frac{1}{2\pi}\int_{\mathcal{C}_{\epsilon}}|z^{1/2}|\cdot\left|\operatorname{det}\left(h_{z}\left(H_{k}\right)\right)\right|\cdot\left\|h_{z}(M)\right\|\,|\mathrm{d}z|\right)\left\|\bm{\xi}_{0}^{k}\right\|$
-$+\left(\frac{1}{2\pi}\int_{\mathcal{C}_{\delta}^{+}}|z^{1/2}|\cdot\left|\operatorname{det}\left(h_{z}\left(H_{k}\right)\right)\right|\cdot\left\|h_{z}(M)\right\|\,|\mathrm{d}z|\right)\left\|\bm{\xi}_{0}^{k}\right\|$
-(9) $-\left(\frac{1}{2\pi}\int_{\mathcal{C}_{\delta}^{-}}|z^{1/2}|\cdot\left|\operatorname{det}\left(h_{z}\left(H_{k}\right)\right)\right|\cdot\left\|h_{z}(M)\right\|\,|\mathrm{d}z|\right)\left\|\bm{\xi}_{0}^{k}\right\|.$
-Let $\mathcal{I}_{1},\ldots,\mathcal{I}_{k}\subset\mathbb{C}$ be subsets such that the eigenvalues of $H_{k}$ are contained in $\mathcal{I}_{i}$ for $i=1,\ldots,k$.
-Define the norm $\|h\|_{\mathcal{I}}:=\sup_{x\in\mathcal{I}}|h(x)|$ for a function $h(x)$ on $\mathcal{I}\subset\mathbb{C}$, where $|h(x)|$ denotes the modulus of $h(x)$.
-Note that the Jordan block definition of matrix functions Eq.
-5 implies the eigenvalues of $f(M)$ are exactly $f(\lambda_{i})$, where $\lambda_{i}$ are the eigenvalues of $M$.
-Thus,
-$\left|\operatorname{det}\left(h_{z}\left(H_{k}\right)\right)\right|=\left|\prod_{i=1}^{k}\lambda_{i}\left(h_{z}\left(H_{k}\right)\right)\right|=\left|\prod_{i=1}^{k}h_{z}\left(\lambda_{i}\left(H_{k}\right)\right)\right|\leq\prod_{i=1}^{k}\left\|h_{z}\right\|_{\mathcal{I}_{i}}.$
-Plugging this into (3.9) yields
-$\left\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(f;M,\bm{b})\right\|\leq$ $\left(\frac{1}{2\pi}\int_{\mathcal{C}_{R}}|z^{1/2}|\cdot\prod_{i=1}^{k}\left\|h_{z}\right\|_{\mathcal{I}_{i}}\cdot\left\|h_{z}(M)\right\|\,|\mathrm{d}z|\right)\left\|\bm{\xi}_{0}^{k}\right\|$
-$-\left(\frac{1}{2\pi}\int_{\mathcal{C}_{\epsilon}}|z^{1/2}|\cdot\prod_{i=1}^{k}\left\|h_{z}\right\|_{\mathcal{I}_{i}}\cdot\left\|h_{z}(M)\right\|\,|\mathrm{d}z|\right)\left\|\bm{\xi}_{0}^{k}\right\|$
-$+\left(\frac{1}{2\pi}\int_{\mathcal{C}_{\delta}^{+}}|z^{1/2}|\cdot\prod_{i=1}^{k}\left\|h_{z}\right\|_{\mathcal{I}_{i}}\cdot\left\|h_{z}(M)\right\|\,|\mathrm{d}z|\right)\left\|\bm{\xi}_{0}^{k}\right\|$
-(10) $-\left(\frac{1}{2\pi}\int_{\mathcal{C}_{\delta}^{-}}|z^{1/2}|\cdot\prod_{i=1}^{k}\left\|h_{z}\right\|_{\mathcal{I}_{i}}\cdot\left\|h_{z}(M)\right\|\,|\mathrm{d}z|\right)\left\|\bm{\xi}_{0}^{k}\right\|.$
-As the outer radius $R\to\infty$, the integral over the large circular arc vanishes for $k\geq 1$ since $|z^{1/2}|=O\left(R^{1/2}\right)$, $\left\|h_{z}\right\|_{\mathcal{I}_{i}}=O\left(R^{-1}\right)$, and $\|h_{z}(M)\|=O(R^{-1})$.
+$\left\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(f;M,\bm{b})\right\|$
+$\leq$ $\left(\frac{1}{2\pi}\int_{\mathcal{C}_{R}}|z^{1/2}|\prod_{i=1}^{k}\left|\frac{\lambda_{i}\left(H_{k}\right)}{\lambda_{i}(H_{k})-z}\right|\left\|(M-zI)^{-1}M\right\|\,|\mathrm{d}z|\right)\left\|\bm{\xi}_{0}^{k}\right\|$
+$-\left(\frac{1}{2\pi}\int_{\mathcal{C}_{\epsilon}}|z^{1/2}|\prod_{i=1}^{k}\left|\frac{\lambda_{i}\left(H_{k}\right)}{\lambda_{i}(H_{k})-z}\right|\left\|(M-zI)^{-1}M\right\|\,|\mathrm{d}z|\right)\left\|\bm{\xi}_{0}^{k}\right\|$
+$+\left(\frac{1}{2\pi}\int_{\mathcal{C}_{\delta}^{+}}|z^{1/2}|\prod_{i=1}^{k}\left|\frac{\lambda_{i}\left(H_{k}\right)}{\lambda_{i}(H_{k})-z}\right|\left\|(M-zI)^{-1}M\right\|\,|\mathrm{d}z|\right)\left\|\bm{\xi}_{0}^{k}\right\|$
+$-\left(\frac{1}{2\pi}\int_{\mathcal{C}_{\delta}^{-}}|z^{1/2}|\prod_{i=1}^{k}\left|\frac{\lambda_{i}\left(H_{k}\right)}{\lambda_{i}(H_{k})-z}\right|\left\|(M-zI)^{-1}M\right\|\,|\mathrm{d}z|\right)\left\|\bm{\xi}_{0}^{k}\right\|.$
+As the outer radius $R\to\infty$, the integral over the large circular arc vanishes for $k\geq 1$ since $|z^{1/2}|=\mathcal{O}\left(R^{1/2}\right)$, $\left|\frac{\lambda_{i}(H_{k})}{\lambda_{i}(H_{k})-z}\right|=\mathcal{O}(R^{-1})$, and $\|M(M-zI)^{-1}\|=\mathcal{O}(R^{-1})$.
 Similarly, as the inner radius $\epsilon\to 0$, the integral goes to zero.
@@
 $\left\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(f;M,\bm{b})\right\|$
-$\leq$ $\lim_{\begin{subarray}{c}\epsilon\to 0\\ R\to\infty\end{subarray}}\left(\frac{1}{2\pi}\int_{-R}^{-\epsilon}\left|(x+\delta i)^{1/2}\right|\cdot\left(\prod_{i=1}^{k}\left\|h_{x+\delta i}\right\|_{\mathcal{I}_{i}}\right)\cdot\left\|h_{x+\delta i}(H_{k})\right\|\mathrm{~{}d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|$
-$-\left(\frac{1}{2\pi}\int^{-R}_{-\epsilon}\left|(x-\delta i)^{1/2}\right|\cdot\left(\prod_{i=1}^{k}\left\|h_{x-\delta i}\right\|_{\mathcal{I}_{i}}\right)\cdot\left\|h_{x-\delta i}(H_{k})\right\|\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|$
-$=$ $\left(\frac{1}{\pi}\int_{0}^{\infty}x^{1/2}\cdot\left(\prod_{i=1}^{k}\left\|h_{-x}\right\|_{\mathcal{I}_{i}}\right)\cdot\left\|h_{-x}(M)\right\|\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|.$
-Step 2: Bound $\left\|h_{-x}\right\|_{\mathcal{I}_{i}}$ and $\left\|h_{-x}(M)\right\|$. Consider any $w\in\mathbb{C}$ and a fixed $x\in\mathbb{R}$ such that $x>0$, then,
-$|w+x|=\sqrt{(\operatorname{Re}(w)+x)^{2}+\operatorname{Im}(w)^{2}}\quad\text{and}\quad\sqrt{|w|^{2}+x^{2}}=\sqrt{\operatorname{Re}(w)^{2}+\operatorname{Im}(w)^{2}+x^{2}}.$
-A straightforward calculation then yields
-$|h_{-x}(w)|=\frac{|w|}{|w+x|}\leq\frac{|w|}{\sqrt{|w|^{2}+x^{2}}}:=g_{x}(w).$
-The function $g_{x}(w)$ is increasing with respect to $|w|$. Noting that $H_{k}=Q_{k}^{*}MQ_{k}$, we apply the interlacing inequalities for singular values in Theorem 2 of [34] to obtain:
-$\sigma_{\max}(H_{k})\leq\sigma_{\max}(M)\quad\text{and}\quad\sigma_{\min}(H_{k})\geq\sigma_{\min}(M).$
-Furthermore, the modulus of the eigenvalues of a matrix is bounded above by its largest singular value, and the modulus of the smallest eigenvalue is bounded below by its smallest singular value.
-Letting $\mathcal{S}:=[\sigma_{\min}(M),\sigma_{\max}(M)]$, the problem now turns into analyzing the behavior of $g_{x}$ within $\mathcal{S}$,
-(11) $\left\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(f;M,\bm{b})\right\|\leq\left(\frac{1}{\pi}\int_{0}^{\infty}x^{1/2}\cdot\left(\prod_{i=1}^{k}\left\|g_{x}\right\|_{\mathcal{S}}\right)\cdot\left\|h_{-x}(M)\right\|\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|.$
-We next give the bound for $\left\|h_{-x}(M)\right\|$.
-For a fixed real number $x>0$, we consider a new contour $\mathcal{C}$ as in Fig.
-1, but with $R=2\sigma_{\max}(M)$ and $\delta=\frac{\epsilon C_{M}}{\sqrt{C_{M}^{2}+4}}$, where $C_{M}=\frac{\lambda_{\min}\left(\frac{M+M^{*}}{2}\right)}{\left|\lambda_{\min}\left(i\left(\frac{M-M^{*}}{2}\right)\right)\right|}$.
-Using the Cauchy integral formula, it follows that
-$h_{-x}(M)=-\frac{1}{2\pi i}\int_{\mathcal{C}}h_{-x}(s)(M-sI)^{-1}\,\mathrm{d}s.$
-Then, we have:
-$\left\|h_{-x}(M)\right\|$
-$\leq$ $\frac{1}{2\pi}\int_{\mathcal{C}_{R}}|h_{-x}(s)|\cdot\left\|(M-sI)^{-1}\right\|\,|\mathrm{d}s|-\frac{1}{2\pi}\int_{\mathcal{C}_{\epsilon}}|h_{-x}(s)|\cdot\left\|(M-sI)^{-1}\right\|\,|\mathrm{d}s|$
-$+\frac{1}{2\pi}\int_{\mathcal{C}_{\delta}^{+}}|h_{-x}(s)|\cdot\left\|(M-sI)^{-1}\right\|\,|\mathrm{d}s|-\frac{1}{2\pi}\int_{\mathcal{C}_{\delta}^{-}}|h_{-x}(s)|\cdot\left\|(M-sI)^{-1}\right\|\,|\mathrm{d}s|$
-$\leq$ $\frac{1}{2\pi}\int_{\mathcal{C}_{R}}|g_{x}(s)|\cdot\left\|(M-sI)^{-1}\right\|\,|\mathrm{d}s|-\frac{1}{2\pi}\int_{\mathcal{C}_{\epsilon}}|g_{x}(s)|\cdot\left\|(M-sI)^{-1}\right\|\,|\mathrm{d}s|$
-$+\frac{1}{2\pi}\int_{\mathcal{C}_{\delta}^{+}}|g_{x}(s)|\cdot\left\|(M-sI)^{-1}\right\|\,|\mathrm{d}s|-\frac{1}{2\pi}\int_{\mathcal{C}_{\delta}^{-}}|g_{x}(s)|\cdot\left\|(M-sI)^{-1}\right\|\,|\mathrm{d}s|.$
-Since the integral on the small circular arc tends to 00 as $\epsilon\to 0$, we only need to take care of the integral along $\mathcal{C}^{+}_{\delta}$, $\mathcal{C}^{-}_{\delta}$ and over the large circular arc $\mathcal{C}_{R}$.
-On the large circular arc $\mathcal{C}_{R}$, we start with the expression:
-$\left\|(M-sI)^{-1}\right\|$ $=\frac{1}{\sqrt{\lambda_{\min}\left((M-sI)^{*}(M-sI)\right)}}$
-$=\frac{1}{\sqrt{\lambda_{\min}\left(M^{*}M-s^{*}M-sM^{*}+|s|^{2}I\right)}}$
-(12) $\leq\frac{1}{\sqrt{\lambda_{\min}\left(M^{*}M\right)+\lambda_{\min}\left(|s|^{2}I\right)+\lambda_{\min}\left(-s^{*}M-sM^{*}\right)}}.$
-Letting $T=-s^{*}M-sM^{*}$, we get
-$\bm{x}^{*}T\bm{x}$ $=-s^{*}\bm{x}^{*}M\bm{x}-s\bm{x}^{*}M^{*}\bm{x}$
-$=-s^{*}\bm{x}^{*}M\bm{x}-s(\bm{x}^{*}M\bm{x})^{*}.$
-Thus, one obtains $\lambda_{\min}(T)=\min\limits_{\|\bm{x}\|=1}\bm{x}^{*}T\bm{x}=-2\max\limits_{\|\bm{x}\|=1}\operatorname{Re}(s^{*}\bm{x}^{*}M\bm{x}).$ In addition, applying the Cauchy-Schwarz inquality leads to
-$\operatorname{Re}(s^{*}\bm{x}^{*}M\bm{x})\leq|s^{*}\bm{x}^{*}M\bm{x}|=|s|\cdot|\bm{x}^{*}M\bm{x}|\leq|s|\cdot\|M\|=|s|\sigma_{\max}(M).$
-Then, we find $\lambda_{\min}(T)\geq-2|s|\sigma_{\max}(M)$. Substituting the inequality into (3.9), we have
-$\left\|(M-sI)^{-1}\right\|$ $\leq\frac{1}{\sqrt{(\sigma_{\min}(M))^{2}+|s|^{2}-2|s|\sigma_{\max}(M)}}$
-$=\frac{1}{\sigma_{\min}(M)}.$
-Thus, over the the large circular arc $\mathcal{C}_{R}$, we obtain the following estimate:
-$\frac{1}{2\pi}\int_{\mathcal{C}_{R}}|g_{x}(s)|\cdot\left\|(M-sI)^{-1}\right\||\mathrm{d}s|\leq$ $\frac{2\pi R}{2\pi\sigma_{\min}(M)}\frac{R}{\sqrt{R^{2}+x^{2}}}$
-$=$ $\frac{4\sigma_{\max}(M)\kappa(M)}{(\sqrt{4(\sigma_{\max}(M))^{2}+x^{2}})}.$
-Along the lines $\mathcal{C}^{+}_{\delta}$, $\mathcal{C}^{-}_{\delta}$, one arrives at
-$\left\|(M-sI)^{-1}\right\|$ $=\frac{1}{\sqrt{\lambda_{\min}\left(M^{*}M-s^{*}M-sM^{*}+|s|^{2}I\right)}}$
-(13) $\leq\frac{1}{\sqrt{\lambda_{\min}\left(M^{*}M\right)+\lambda_{\min}(-s^{*}M-sM^{*})}}.$
-Given that $M$ is positive-definite, $\frac{M+M^{*}}{2}$ is also positive-definite.
-We assume $s$ takes the form: $s=s_{1}\pm i\delta$.
-Since $\delta=\frac{\epsilon C_{M}}{\sqrt{C_{M}^{2}+4}}$ with $C_{M}=\frac{\lambda_{\min}\left(\frac{M+M^{*}}{2}\right)}{\left|\lambda_{\min}\left(i\left(\frac{M-M^{*}}{2}\right)\right)\right|}$, then $-R\leq s_{1}\leq-\frac{2\epsilon}{\sqrt{C_{M}^{2}+4}}<-\epsilon$.
-Therefore,
-$\lambda_{\min}(-s^{*}M-sM^{*})$ $=-2\lambda_{\max}\left(s_{1}\left(\frac{M+M^{*}}{2}\right)-i\delta\left(\frac{M-M^{*}}{2}\right)\right)$
-$\geq-2\lambda_{\max}\left(s_{1}\left(\frac{M+M^{*}}{2}\right)\right)-2\lambda_{\max}\left(-i\delta\left(\frac{M-M^{*}}{2}\right)\right)$
-$=-2s_{1}\lambda_{\min}\left(\frac{M+M^{*}}{2}\right)+2\delta\lambda_{\min}\left(i\left(\frac{M-M^{*}}{2}\right)\right)$
-$\geq 0.$
-Substituting this into 3.9 yields
-$\left\|(M-sI)^{-1}\right\|\leq\frac{1}{\sigma_{\min}(M)}.$
-Analogously, with $\delta=\frac{\epsilon C_{M}}{\sqrt{C_{M}^{2}+4}}$, we get
-$\frac{1}{2\pi}\int_{\mathcal{C}_{\delta}^{+}}|g_{x}(s)|\cdot\left\|(M-sI)^{-1}\right\|\,|\mathrm{d}s|-\frac{1}{2\pi}\int_{\mathcal{C}_{\delta}^{-}}|g_{x}(s)|\cdot\left\|(M-sI)^{-1}\right\|\,|\mathrm{d}s|$
-$\leq$ $\frac{1}{2\pi}\lim_{\begin{subarray}{c}\epsilon\to 0\end{subarray}}\int_{-R}^{-\epsilon}|g_{x}(s_{1}+\delta i)|\cdot\left\|(M-(s_{1}+\delta i)I)^{-1}\right\|\mathrm{d}s_{1}$
-$-\frac{1}{2\pi}\lim_{\begin{subarray}{c}\epsilon\to 0\end{subarray}}\int^{-R}_{-\epsilon}|g_{x}(s_{1}-\delta i)|\cdot\left\|(M-(s_{1}-\delta i)I)^{-1}\right\|\mathrm{d}s_{1}$
-$=$ $\frac{1}{\pi}\int_{0}^{R}|g_{x}(-s_{1})|\cdot\left\|(M+s_{1}I)^{-1}\right\|\mathrm{d}s_{1}$
-$\leq$ $\frac{R^{2}}{\pi\sigma_{\min}(M)(\sqrt{R^{2}+x^{2}})}$
-$=$ $\frac{4\sigma_{\max}(M)\kappa(M)}{\pi(\sqrt{4(\sigma_{\max}(M))^{2}+x^{2}})}.$
-Finally, combining all estimates, we obtain the upper bound:
-(14) $\left\|h_{-x}(M)\right\|\leq\frac{1+\pi}{\pi}\frac{4\sigma_{\max}(M)\kappa(M)}{(\sqrt{4(\sigma_{\max}(M))^{2}+x^{2}})}\leq\frac{4(1+\pi)\kappa(M)}{\pi}\frac{\sigma_{\max}(M)}{(\sqrt{(\sigma_{\max}(M))^{2}+x^{2}})}.$
-Step 3: Apply Beta and Gamma Functions. Substituting (14) into (11) gives
-$\left\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(f;M,\bm{b})\right\|$
-$\leq\left(\frac{1}{\pi}\int_{0}^{\infty}x^{1/2}\cdot\left\|g_{x}(M)\right\|_{\mathcal{S}}^{k}\cdot\left\|h_{-x}\right\|\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|$
-$\leq\frac{4(1+\pi)\kappa(M)}{\pi^{2}}\left(\int_{0}^{\infty}x^{1/2}\cdot\left\|g_{x}\right\|_{\mathcal{S}}^{k}\cdot\frac{\sigma_{\max}(M)}{(\sqrt{(\sigma_{\max}(M))^{2}+x^{2}})}\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|$
-$\leq\frac{4(1+\pi)\kappa(M)}{\pi^{2}}\left(\int_{0}^{\infty}x^{1/2}\cdot\frac{(\sigma_{\max}(M))^{k+1}}{(\sqrt{(\sigma_{\max}(M))^{2}+x^{2}})^{k+1}}\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|.$
-Define $\mathrm{I}:=\int_{0}^{\infty}x^{1/2}\cdot\frac{(\sigma_{\max}(M))^{k+1}}{(\sqrt{(\sigma_{\max}(M))^{2}+x^{2}})^{k+1}}\,\mathrm{d}x$. Let $x=\sigma_{\max}(M)y$, then we obtain
-$\rm{I}$ $=(\sigma_{\max}(M))^{3/2}\int^{\infty}_{0}\frac{y^{1/2}}{(\sqrt{1+y^{2}})^{k+1}}\mathrm{~{}d}y$
-$=(\sigma_{\max}(M))^{3/2}\int^{\frac{\pi}{2}}_{0}\frac{(\tan(\theta))^{1/2}(\sec(\theta))^{2}}{(\sec(\theta))^{k+1}}\mathrm{~{}d}\theta$
-$=(\sigma_{\max}(M))^{3/2}\int^{\frac{\pi}{2}}_{0}(\sin(\theta))^{1/2}(\cos(\theta))^{k-3/2}\mathrm{~{}d}\theta$
-$=\frac{(\sigma_{\max}(M))^{3/2}}{2}\mathrm{B}\left(\frac{3}{4},\frac{k}{2}-\frac{1}{4}\right)$
-$=\frac{(\sigma_{\max}(M))^{3/2}}{2}\frac{\Gamma(3/4)\Gamma(k/2-1/4)}{\Gamma(k/2+1/2)}.$
-Here, $\text{B}(\cdot,\cdot)$ and $\Gamma(\cdot)$ are the standard Beta and Gamma functions, respectively, with $\mathrm{B}(x,y)=\frac{\Gamma(x)\Gamma(y)}{\Gamma(x+y)}.$ Furthermore, noting that in [36] for $w>0$ and $l\in(0,1)$,
-$\left(\frac{w}{w+l}\right)^{1-l}\leq\frac{\Gamma(w+l)}{w^{l}\Gamma(w)}\leq 1,$
-we arrive at
-$\frac{\Gamma(k/2-1/4)}{\Gamma(k/2+1/2)}$ $\leq\frac{(k/2+1/2)^{1/4}}{(k/2-1/4)^{1/4}(k/2-1/4)^{3/4}}$
-$\leq\frac{2^{5/4}}{(k-1/2)^{3/4}}$
-$\leq\frac{\sqrt{2}\pi^{2}}{(1+\pi)\Gamma(3/4)(k-1/2)^{3/4}}.$
-Then, it holds that
-$\mathrm{I}\leq\frac{\pi^{2}(\sigma_{\max}(M))^{3/2}}{\sqrt{2}(1+\pi)(k-1/2)^{3/4}}.$
-Collecting all estimates leads to
-$\left\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(f;M,\bm{b})\right\|\leq\frac{2\sqrt{2}\kappa(M)(\sigma_{\max}(M))^{3/2}}{(k-1/2)^{3/4}}\left\|\bm{\xi}_{0}^{k}\right\|.$
-Thus, the relative error of the Arnoldi approximation for matrix square roots is bounded by:
-$\frac{\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};M,\bm{b})\|}{\|M^{1/2}\bm{b}\|}$ $\leq\frac{2\sqrt{2}\kappa(M)(\sigma_{\max}(M))^{3/2}}{(k-1/2)^{3/4}}\frac{\left\|\bm{\xi}_{0}^{k}\right\|}{\|M^{1/2}\bm{b}\|}$
-(15) $\leq\frac{2\sqrt{2}(\kappa(M))^{5/2}}{(k-1/2)^{3/4}}\frac{\|\bm{r}^{k}_{0}\|}{\|\bm{b}\|}.$
-Step 4: Bound $\|\bm{r}^{k}_{0}\|$. From (6), it holds that
-$\bm{r}^{k}_{0}$ $=\bm{b}-MQ_{k}H_{k}^{-1}\|\bm{b}\|\bm{e}_{1}$
-$=\bm{b}-(Q_{k}H_{k}+h_{k+1,k}\bm{q}_{k+1}\bm{e}_{k}^{T})H_{k}^{-1}\|\bm{b}\|\bm{e}_{1}$
-$=-h_{k+1,k}\bm{q}_{k+1}\bm{e}_{k}^{T}H_{k}^{-1}\|\bm{b}\|\bm{e}_{1}.$
-Taking the 2-norm on both sides gives
-(16) $\left\|\bm{r}^{k}_{0}\right\|\leq h_{k+1,k}\left\|H_{k}^{-1}\right\|\left\|\bm{b}\right\|\leq\frac{h_{k+1,k}}{\sigma_{\min}\left(H_{k}\right)}\left\|\bm{b}\right\|.$
-Multiplying $\bm{e}_{k}$ on both sides of (6) and then taking the 2-norm, we have
-(17) $h_{k+1,k}=\|M\bm{q}_{k}-Q_{k}H_{k}\bm{e}_{k}\|\leq\|M\bm{q}_{k}\|+\|H_{k}\bm{e}_{k}\|\leq\sigma_{\max}(M)+\sigma_{\max}(H_{k}),$
-where the second inequality is achieved by utilizing the triangle inequality and the orthogonality of $Q_{k}$.
-Plugging the above estimate into (16) and utilizing the interlacing inequalities for singular values, we further bound $\left\|\bm{r}_{0}^{k}\right\|$ by
-$\left\|\bm{r}^{k}_{0}\right\|\leq\frac{\sigma_{\max}(M)+\sigma_{\max}(H_{k})}{\sigma_{\min}\left(H_{k}\right)}\left\|\bm{b}\right\|\leq\frac{2\sigma_{\max}\left(M\right)}{\sigma_{\min}\left(M\right)}\left\|\bm{b}\right\|=2\kappa(M)\|\bm{b}\|.$
-Consequently, substituting the above estimate of $\bm{r}_{0}^{k}$ into (3.9), we obtain
-$\frac{\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};M,\bm{b})\|}{\|M^{1/2}\bm{b}\|}\leq\frac{2\sqrt{2}(\kappa(M))^{5/2}}{(k-1/2)^{3/4}}\frac{\|\bm{r}^{k}_{0}\|}{\|\bm{b}\|}\leq\frac{4\sqrt{2}(\kappa(M))^{7/2}}{(k-1/2)^{3/4}}.$
+$\leq$ $\lim_{\begin{subarray}{c}\epsilon\to 0\\ R\to\infty\end{subarray}}\left(\frac{1}{2\pi}\int_{-R}^{-\epsilon}\left|(x+\delta i)^{1/2}\right|\prod_{i=1}^{k}\left|\frac{\lambda_{i}(H_{k})}{\lambda_{i}(H_{k})-(x+\delta_{i})}\right|\left\|\left(M-(x+\delta iI)\right)^{-1}M\right\|\mathrm{~d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|$
+$-\left(\frac{1}{2\pi}\int^{-R}_{-\epsilon}\left|(x-\delta i)^{1/2}\right|\prod_{i=1}^{k}\left|\frac{\lambda_{i}(H_{k})}{\lambda_{i}(H_{k})-(x+\delta_{i})}\right|\left\|\left(M-(x-\delta iI)\right)^{-1}M\right\|\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|$
+$=$ $\frac{1}{\pi}\left(\int_{0}^{\infty}x^{1/2}\prod_{i=1}^{k}\left|\frac{\lambda_{i}(H_{k})}{\lambda_{i}(H_{k})+x}\right|\left\|(M+xI)^{-1}M\right\|\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|.$
+Next, we prove that $\left\|(M+xI)^{-1}M\right\|<1$. This is equivalent to showing that for any non-zero vector $v\in\mathbb{C}^{n}$ and $x>0$,
+$\left\|(M+xI)^{-1}M\bm{v}\right\|<\left\|\bm{v}\right\|.$
+To prove this, let $\bm{u}=(M+xI)^{-1}M\bm{v}$, and note that $M(\bm{u}-\bm{v})=-x\bm{u}.$ Taking the inner product with $\bm{u}-\bm{v}$, we have
+$x\operatorname{Re}\left(\bm{u},\bm{u}-\bm{v}\right)=\operatorname{Re}\left(M(\bm{u}-\bm{v}),\bm{u}-\bm{v}\right)>0.$
+Thus, we have
+$0>\operatorname{Re}\left(\bm{u},\bm{u}-\bm{v}\right)=\left\|\bm{u}\right\|^{2}-\operatorname{Re}\left(\bm{u},\bm{v}\right)\geq\left\|\bm{u}\right\|^{2}-\|\left(\bm{u},\bm{v}\right)\|\geq\left\|\bm{u}\right\|^{2}-\left\|\bm{u}\right\|\left\|\bm{v}\right\|$
+which implies $\left\|\bm{u}\right\|<\left\|\bm{v}\right\|$ and, therefore, $\left\|(M+xI)^{-1}M\right\|<1$. Thus,
+$\left\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(f;M,\bm{b})\right\|=\frac{1}{\pi}\left(\int_{0}^{\infty}x^{1/2}\prod_{i=1}^{k}\left|\frac{\lambda_{i}(H_{k})}{\lambda_{i}(H_{k})+x}\right|\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|,$
+which completes the proof.
 Remark 3.10.
-If the matrix $M$ is Hermitian, then using the results from [6] and [1], we obtain
-$\frac{\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};M,\bm{b})\|}{\|M^{1/2}\bm{b}\|}\leq\frac{(\kappa(M))^{3/2}}{2k^{3/2}}\frac{\|\bm{r}^{k}_{0}\|}{\|\bm{b}\|}\leq\frac{(\kappa(M))^{5/2}}{k^{3/2}}.$
-This sharper result stems from the fact that the Hermitian matrix is unitarily diagonalizable with real eigenvalues, which allows for a more refined spectral analysis of $\left|\operatorname{det}\left(h_{z}\left(H_{k}\right)\right)\right|$ and $\left\|h_{z}(M)\right\|$ in 3.9.
+In the Arnoldi process, $H_{k}$ is generated naturally, and the FOM error $\left\|\bm{\xi}_{0}^{k}\right\|$ can be obtained with essentially one additional step.
+Therefore, the error bound in Theorem 3.8 provides an a posteriori error estimate that can be computed efficiently in practice, especially for small $k$.
+This motivates the following practical stopping criterion:
+$\frac{1}{\pi}\left\|\bm{\xi}_{0}^{k}\right\|\int_{0}^{\infty}x^{1/2}\prod_{i=1}^{k}\left|\frac{\lambda_{i}(H_{k})}{\lambda_{i}(H_{k})+x}\right|\,\mathrm{d}x\leq\textit{tol}.$
+where tol is a given tolerance.
+Next, building on the a posteriori error estimate in Theorem 3.8, we derive an a priori error estimate, which is summarized in the following theorem.
+Theorem 3.11.
+Given a positive-definite matrix $M\in\mathbb{C}^{n\times n}$ and a vector $\bm{b}\in\mathbb{C}^{n}$, for $k\geq 2$, the error of the Arnoldi process for approximating the matrix square root is bounded by
+$\left\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(f;M,\bm{b})\right\|\leq\frac{\Gamma\!(3/4)}{2^{1/4}\pi}\frac{2k}{2k-3}(\sigma_{\max}(M))^{\frac{3}{2}}k^{-\frac{3}{4}}\left\|\bm{\xi}_{0}^{k}\right\|,$
+where $\Gamma\!(\cdot)$ is the standard Gamma function, and $\sigma_{\max}(M)$ denotes the largest singular value of $M$.
+Proof 3.12.
+From Theorem 3.8, we know that
+$\left\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(f;M,\bm{b})\right\|\leq\frac{1}{\pi}\int_{0}^{\infty}x^{1/2}\prod_{i=1}^{k}\left|\frac{\lambda_{i}(H_{k})}{\lambda_{i}(H_{k})+x}\right|\,\mathrm{d}x.$
+Consider any $w\in\mathbb{C}$ with $\operatorname{Re}(w)>0$ and fix $x>0$ in $\mathbb{R}$, a straightforward calculation then yields
+$\left|\frac{w}{w+x}\right|\leq\frac{|w|}{\sqrt{|w|^{2}+x^{2}}}.$
+The function on the right-hand side is increasing with respect to $|w|$.
+Denote $\sigma_{\min}(M)$ as the smallest eigenvalue of $M$.
+Applying the interlacing inequalities for singular values in Theorem 2 of [thompson1972principal] gives
+$\max_{i}|\lambda_{i}(H_{k})|\leq\sigma_{\max}(H_{k})\leq\sigma_{\max}(M),\qquad\min_{i}|\lambda_{i}(H_{k})|\geq\sigma_{\min}(H_{k})\geq\sigma_{\min}(M).$
+Letting $\mathcal{S}:=[\sigma_{\min}(M),\sigma_{\max}(M)]$, one arrives at
+$\int_{0}^{\infty}x^{1/2}\prod_{i=1}^{k}\left|\frac{\lambda_{i}(H_{k})}{\lambda_{i}(H_{k})+x}\right|\,\mathrm{d}x$ $\leq\int_{0}^{\infty}x^{1/2}\prod_{i=1}^{k}\frac{|\lambda_{i}(H_{k})|}{\sqrt{|\lambda_{i}(H_{k})|^{2}+x^{2}}}\,\mathrm{d}x$
+$\leq\int_{0}^{\infty}x^{1/2}\left(\sup\limits_{|w|\in\mathcal{S}}\frac{|w|}{\sqrt{|w|^{2}+x^{2}}}\right)^{k}\,\mathrm{d}x$
+$\leq\int_{0}^{\infty}x^{1/2}\frac{(\sigma_{\max}(M))^{k}}{(\sqrt{(\sigma_{\max}(M))^{2}+x^{2}})^{k}}\,\mathrm{d}x.$
+Letting $x=\sigma_{\max}(M)\tan(\theta)$, for $k\geq 2$, we obtain
+$\int_{0}^{\infty}x^{1/2}\frac{(\sigma_{\max}(M))^{k}}{\bigl(\sqrt{(\sigma_{\max}(M))^{2}+x^{2}}\bigr)^{k}}\,\mathrm{d}x$ $=(\sigma_{\max}(M))^{3/2}\int_{0}^{\frac{\pi}{2}}\frac{\tan^{1/2}(\theta)\,\sec^{2}(\theta)}{\sec^{k}(\theta)}\,\mathrm{d}\theta$
+$=(\sigma_{\max}(M))^{3/2}\int_{0}^{\frac{\pi}{2}}\sin^{1/2}(\theta)\,\cos^{k-5/2}(\theta)\,\mathrm{d}\theta$
+$=\frac{(\sigma_{\max}(M))^{3/2}}{2}\,\mathrm{B}\!\left(\frac{3}{4},\,\frac{2k-3}{4}\right)$
+$=\frac{(\sigma_{\max}(M))^{3/2}}{2}\,\frac{\Gamma\!\left(3/4\right)\Gamma\!\left((2k-3)/4\right)}{\Gamma\!\left(k/2\right)},$
+where $\mathrm{B}(\cdot,\cdot)$ is the Beta function, satisfying $\mathrm{B}(x,y)=\frac{\Gamma(x)\Gamma(y)}{\Gamma(x+y)}$. Furthermore, following [wendel1948note], for $w>0$ and $l\in(0,1)$,
+$\frac{w}{(w+l)^{1-l}}\leq\frac{\Gamma(w+l)}{\Gamma(w)}\leq w^{l},$
+Then, with w=(2k3)/4w=(2k-3)/4 and $l=3/4$, we have,
+$\frac{\Gamma\!\left((2k-3)/4\right)}{\Gamma\!\left(k/2\right)}\leq\frac{4}{2k-3}\left(\frac{k}{2}\right)^{\frac{1}{4}}=2^{\frac{7}{4}}\frac{k^{1/4}}{2k-3},$
+Collecting all estimates yields the desired result.
 Remark 3.13.
-Note that the residual $\bm{r}_{0}^{k}$ is precisely the residual from the FOM applied to the linear system $M\bm{x}=\bm{b}$.
-Since the error in Theorem 3.8 is bounded in terms of $\bm{r}_{0}^{k}$, which is naturally computed in the Arnoldi iteration, we thus have a practical stopping criterion for our algorithm.
+Summarizing the above results, we obtain the following estimates:
+(9) $\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};M,\bm{b})\|$ $\leq\frac{1}{\pi}\left(\int_{0}^{\infty}x^{1/2}\prod_{i=1}^{k}\left|\frac{\lambda_{i}(H_{k})}{\lambda_{i}(H_{k})+x}\right|\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|$
+(10) $\leq\frac{1}{\pi}\left(\int_{0}^{\infty}x^{1/2}\prod_{i=1}^{k}\frac{|\lambda_{i}(H_{k})|}{\sqrt{|\lambda_{i}(H_{k})|^{2}+x^{2}}}\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|$
+(11) $\leq\frac{\Gamma\!(3/4)}{2^{1/4}\pi}\frac{2k}{2k-3}(\sigma_{\max}(M))^{\frac{3}{2}}k^{-\frac{3}{4}}\left\|\bm{\xi}_{0}^{k}\right\|.$
+The inequalities Eq.
+9 and Eq.
+10 provide a posteriori error bounds and remain relatively tight.
+The last inequality Eq.
+11 gives an a priori bound; since it relies on $|\lambda_{i}(H_{k})|\leq\sigma_{\max}(M)$, it becomes loose when the eigenvalues of $M$ are widely distributed.
 Remark 3.14.
-The bound in (17) is a loose upper bound for $h_{k+1,k}$.
-In practice, $h_{k+1,k}$ tends to decrease with $k$, as the Krylov subspace increasingly approximates the spectral subspace associated with the dominant eigenvalues of $M$.
-However, this behavior is not guaranteed in general.
-To further understand the behavior of $\|\bm{r}_{0}^{k}\|$, one may explore results from [7].
-They show that $\min_{0\leq j\leq k}\|\mathbf{r}_{0}^{j}\|\leq\sqrt{k+1}\cdot\|\mathbf{r}^{k}_{\mathrm{G}}\|$, where $\mathbf{r}^{k}_{\mathrm{G}}$ denotes the GMRES residual after $k$ iterations for solving $M\bm{x}=\bm{b}$.
-This implies that the overall convergence of FOM is at most a factor of $\sqrt{k+1}$ worse than that of GMRES.
-However, the residual norm at a fixed iteration $k$ can be arbitrarily large, and thus this result does not apply to our setting.
+If the matrix $M$ is Hermitian, then using the results from [chen2022error] and [amsel2023near], we obtain
+$\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};M,\bm{b})\|$
+$\leq\frac{1}{\pi}\left(\int_{0}^{\infty}x^{1/2}\prod_{i=1}^{k}\frac{\lambda_{i}(H_{k})}{\lambda_{i}(H_{k})+x}\frac{\lambda_{\max}(M)}{\lambda_{\max}(M)+x}\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|$
+$\leq\frac{1}{\pi}\left(\int_{0}^{\infty}x^{1/2}\left(\frac{\lambda_{\max}(M)}{\lambda_{\max}(M)+x}\right)^{k+1}\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|$
+(12) $\leq\frac{1}{2\sqrt{\pi}}\frac{2k}{2k-1}(\lambda_{\max}(M))^{\frac{3}{2}}k^{-\frac{3}{2}}\|\bm{\xi}^{k}_{0}\|.$
+This sharper result stems from the fact that the Hermitian matrix is unitarily diagonalizable with real eigenvalues, which allows for a more refined spectral analysis of $\left|\frac{\det(H_{k})}{\det(H_{k}-zI)}\right|$ and $\|(M+xI)^{-1}M\|$.
+Thus, as a byproduct of our analysis, we derive an error bound for the Hermitian case that is tighter than (12), the result from [chen2022error].
+Theorem 3.15.
+Given a positive-definite Hermitian matrix $M\in\mathbb{C}^{n\times n}$ and a vector $\bm{b}\in\mathbb{C}^{n}$, the error of the Arnoldi process for approximating the matrix square root is bounded by
+(13) $\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};M,\bm{b})\|\leq\frac{1}{2\sqrt{\pi}}\frac{2k}{2k-1}\left(\bar{\lambda}_{k+1}(M)\right)^{\frac{3}{2}}k^{-\frac{3}{2}}\left\|\bm{\xi}_{0}^{k}\right\|,$
+where
+$\bar{\lambda}_{k+1}(M):=\frac{1}{k+1}\left(\sum_{i=1}^{k}\lambda_{i}(M)+\lambda_{1}(M)\right)\leq\lambda_{\max}(M).$
+Proof 3.16.
+Let $\lambda_{1}(M)\geq\lambda_{2}(M)\geq\cdots\geq\lambda_{n}(M)$ be the eigenvalues of $M$. Then, $\lambda_{\max}(M):=\lambda_{1}(M)$. Using the arithmetic–geometric mean inequality, we obtain
+$\prod_{i=1}^{k}\frac{\lambda_{i}(H_{k})}{\lambda_{i}(H_{k})+x}\frac{\lambda_{1}(M)}{\lambda_{1}(M)+x}\leq\left(\frac{1}{k+1}\left(\sum_{i=1}^{k}\frac{\lambda_{i}(H_{k})}{\lambda_{i}(H_{k})+x}+\frac{\lambda_{1}(M)}{\lambda_{1}(M)+x}\right)\right)^{k+1}.$
+Observing that $g_{x}(\lambda):=\frac{\lambda}{\lambda+x}$ is a concave function in $\lambda$, Jensen’s inequality yields
+$\frac{1}{k+1}\left(\sum_{i=1}^{k}g_{x}(\lambda_{i}(H_{k}))+g_{x}(\lambda_{1}(M))\right)\leq g_{x}\left(\frac{\sum_{i=1}^{k}\lambda_{i}(H_{k})+\lambda_{1}(M)}{k+1}\right).$
+Then,
+$\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};M,\bm{b})\|$
+$\leq\frac{1}{\pi}\left(\int_{0}^{\infty}x^{1/2}\prod_{i=1}^{k}\frac{\lambda_{i}(H_{k})}{\lambda_{i}(H_{k})+x}\frac{\lambda_{1}(M)}{\lambda_{1}(M)+x}\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|$
+$\leq\frac{1}{\pi}\left(\int_{0}^{\infty}x^{1/2}\left(\frac{\frac{1}{k+1}\left(\sum_{i=1}^{k}\lambda_{i}(H_{k})+\lambda_{1}(M)\right)}{\frac{1}{k+1}\left(\sum_{i=1}^{k}\lambda_{i}(H_{k})+\lambda_{1}(M)\right)+x}\right)^{k+1}\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|.$
+Utilizing the interlacing inequality gives
+$\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};M,\bm{b})\|$ $\leq\frac{1}{\pi}\left(\int_{0}^{\infty}x^{1/2}\left(\frac{\bar{\lambda}_{k+1}(M)}{\bar{\lambda}_{k+1}(M)+x}\right)^{k+1}\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|$
+$\leq\frac{1}{2\sqrt{\pi}}\frac{2k}{2k-1}\left(\bar{\lambda}_{k+1}(M)\right)^{\frac{3}{2}}k^{-\frac{3}{2}}\left\|\bm{\xi}_{0}^{k}\right\|.$
+This implies the bound (13) is sharper then the bound (12) almost for all cases.
+Remark 3.17.
+Additionally, one can similarly derive an error estimate for the Arnoldi approximation of the matrix inverse square root:
+$\|M^{-1/2}\bm{b}-\operatorname{Arn}_{k}(x^{-1/2};M,\bm{b})\|\leq\frac{\Gamma(1/4)}{2^{3/4}\pi}\frac{2k}{2k-1}\left(\sigma_{\max}(M)\right)^{\frac{1}{2}}k^{-\frac{1}{4}}\|\bm{\xi}^{k}_{0}\|.$
+If $M$ is Hermitian, then following [amsel2023near] and the discussion above, we can derive the following tighter bound, which also improves upon the state-of-the-art result,
+$\|M^{-1/2}\bm{b}-\operatorname{Arn}_{k}(x^{-1/2};M,\bm{b})\|\leq\frac{1}{\sqrt{\pi}}\frac{2k+2}{2k+1}\left(\bar{\lambda}_{k+1}(M)\right)^{\frac{1}{2}}(k+1)^{-\frac{1}{2}}\|\bm{\xi}^{k}_{0}\|.$
 3.3 Preprocessing
 In practice, matrices arising from large-scale problems are often too large to be directly measured, stored, or analyzed, necessitating efficient preprocessing.
-To further reduce the cost, one direction is to employ data-sparse approximations, such as hierarchical or low-rank representations [5, 8, 10, 17, 22, 23].
+To further reduce the cost, one direction is to employ data-sparse approximations, such as hierarchical or low-rank representations [bjarkason2019pass, clarkson2017low, datta2010numerical, hackbusch2015hierarchical, kishore2017literature, lin2011fast].
 These methods accelerate matrix-vector multiplications by compressing the underlying structure of the matrix.
-Under this framework, we modify the results from Theorem 3.8 to get an error bound for approximating the square root of a preprocessed matrix using the Arnoldi iteration.
-This is summarized in the following theorem.
-Theorem 3.13.
-Given a positive-definite matrix $M\in\mathbb{C}^{n\times n}$ and a vector $\bm{b}\in\mathbb{C}^{n}$, assume that M admits a data-sparse approximation $\widetilde{M}$ which is positive definite and satisfies, for some $0<\epsilon<1/\kappa(M)$, $\|M-\widetilde{M}\|_{*}\leq\epsilon\|M\|$.
-Then, the relative error of the Arnoldi process for approximating the matrix square root is bounded by
-$\frac{\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};\widetilde{M},\bm{b})\|}{\|M^{1/2}\bm{b}\|}\leq\epsilon^{1/2}\kappa^{1/2}(M)+\frac{4\sqrt{2}(1+\epsilon)^{7/2}(\kappa(M))^{9/2}}{(k-1/2)^{3/4}(1-\epsilon\kappa(M))^{9/2}}.$
-Proof 3.14.
-Based on $\widetilde{M}$, we apply the Arnoldi process to compute the matrix square root. Now the problem turns into:
-$\frac{\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};\widetilde{M},\bm{b})\|}{\|M^{1/2}\bm{b}\|}\leq\frac{\|M^{1/2}\bm{b}-\widetilde{M}^{1/2}\bm{b}\|}{\|M^{1/2}\bm{b}\|}+\frac{\|\widetilde{M}^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};\widetilde{M},\bm{b})\|}{\|M^{1/2}\bm{b}\|}.$
-For the first term on the right hand side, according to the Powers–Størmer inequality [32], we have
-$\frac{\|M^{1/2}\bm{b}-\widetilde{M}^{1/2}\bm{b}\|}{\|M^{1/2}\bm{b}\|}\leq\frac{\|M^{1/2}-\widetilde{M}^{1/2}\|\|\bm{b}\|}{\|M^{1/2}\bm{b}\|}\leq\frac{\|M-\widetilde{M}\|_{*}^{1/2}\|\bm{b}\|}{\|M^{1/2}\bm{b}\|}\leq\epsilon^{1/2}\kappa^{1/2}(M).$
-With Theorem 3.8, the second term is bounded by:
-$\frac{\|\widetilde{M}^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};\widetilde{M},\bm{b})\|}{\|\widetilde{M}^{1/2}\bm{b}\|}\leq\frac{4\sqrt{2}(\kappa(\widetilde{M}))^{7/2}}{(k-1/2)^{3/4}}.$
-By Weyl’s inequality [4], we have $|\sigma_{k}(M)-\sigma_{k}(\widetilde{M})|\leq\|M-\widetilde{M}\|\leq\|M-\widetilde{M}\|_{*}\leq\epsilon\|M\|=\epsilon\sigma_{\max}(M)$. Rearranging terms gives
-$\kappa(\widetilde{M})\leq\frac{(1+\epsilon)\sigma_{\max}(M)}{\sigma_{\min}(M)-\epsilon\sigma_{\max}(M)}=\frac{(1+\epsilon)\kappa(M)}{1-\epsilon\kappa(M)}.$
-Therefore, we conclude that
-$\frac{\|\widetilde{M}^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};\widetilde{M},\bm{b})\|}{\|M^{1/2}\bm{b}\|}$ $\leq\frac{4\sqrt{2}(\kappa(\widetilde{M}))^{7/2}}{(k-1/2)^{3/4}}\frac{\|M^{1/2}\bm{b}\|}{\|\widetilde{M}^{1/2}\bm{b}\|}$
-$\leq\frac{4\sqrt{2}(1+\epsilon)^{7/2}(\kappa(\widetilde{M}))^{9/2}}{(k-1/2)^{3/4}(1-\epsilon\kappa(M))^{9/2}}.$
+Thus, we generalize the results from Theorem 3.8 to get an error bound for approximating the square root of a perturbed matrix using the Arnoldi iteration.
+Corollary 3.18.
+Let $M\in\mathbb{C}^{n\times n}$ be a positive-definite matrix such that for any unit vector $\bm{x}$, $\operatorname{Re}(\bm{x}^{*}M\bm{x})\geq\mu_{1}^{2}$ for some $\mu_{1}>0$, and let $\bm{b}\in\mathbb{C}^{n}$ be a given vector.
+We consider a perturbed matrix $\widetilde{M}$, satisfying $\operatorname{Re}(\bm{x}^{*}\widetilde{M}\bm{x})\geq\mu_{2}^{2}$ for some $\mu_{2}>0$, as well as $\|M-\widetilde{M}\|\leq\epsilon\|M\|$ for some $\epsilon>0$.
+Then, the error of the Arnoldi process for approximating the matrix square root is bounded by
+$\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};\widetilde{M},\bm{b})\|$
+(14) $\leq\frac{\epsilon\sigma_{\max}(M)}{\mu_{1}+\mu_{2}}\|\bm{b}\|+\frac{\Gamma\!(3/4)(1+\epsilon)^{3/2}}{2^{1/4}\pi}\frac{2k}{2k-3}(\sigma_{\max}(M))^{\frac{3}{2}}k^{-\frac{3}{4}}\left\|\bm{\xi}_{0}^{k}\right\|.$
+Proof 3.19.
+Based on $\widetilde{M}$, we apply the Arnoldi process to compute the matrix square root. Now the problem becomes
+$\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};\widetilde{M},\bm{b})\|\leq\|M^{1/2}\bm{b}-\widetilde{M}^{1/2}\bm{b}\|+\|\widetilde{M}^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};\widetilde{M},\bm{b})\|.$
+The first term can be bounded using the perturbation estimate in [schmitt1992perturbation], thus,
+$\|M^{1/2}\bm{b}-\widetilde{M}^{1/2}\bm{b}\|\leq\|M^{1/2}-\widetilde{M}^{1/2}\|\|\bm{b}\|\leq\frac{1}{\mu_{1}+\mu_{2}}\|M-\widetilde{M}\|\|\bm{b}\|\leq\frac{\epsilon\sigma_{\max}(M)}{\mu_{1}+\mu_{2}}\|\bm{b}\|.$
+By Theorem 3.11, the second term is bounded by:
+$\|\widetilde{M}^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};\widetilde{M},\bm{b})\|\leq\frac{\Gamma\!(3/4)}{2^{1/4}\pi}\frac{2k}{2k-3}(\sigma_{\max}(\widetilde{M}))^{\frac{3}{2}}k^{-\frac{3}{4}}\left\|\bm{\xi}_{0}^{k}\right\|.$
+By Weyl’s inequality [bhatia1997matrix], we have $|\sigma_{k}(M)-\sigma_{k}(\widetilde{M})|\leq\|M-\widetilde{M}\|\leq\epsilon\|M\|=\epsilon\sigma_{\max}(M)$.
+Combining the above estimates yields (3.18)
 4 Numerical experiments and discussion
 In this section, we report on numerical studies of the Arnoldi iteration for computing the matrix square root applied to a vector, that is, $M^{1/2}\bm{b}$.
-First, we explore various matrix settings to study the dependence of relative error on the condition number of matrices and the iteration count.
-We start with a Hermitian matrix, confirming the result in Remark 3.10 as well as Lemma 3.3 in [1].
-Then, we test both sparse and general non-Hermitian matrices validating Theorem 3.8.
+We first evaluate the a posteriori bounds (9) and (10), and the a priori bound (11) using non-Hermitian matrices.
+Next, we compare the a priori bounds (12) and (13) in the Hermitian setting to demonstrate the improvement yielded by our new bound (13).
+Then, we investigate the dependence of the approximation error on the largest singular value $\sigma_{\max}(M)$ and the iteration count $k$.
 Finally, we explore the use of the Arnoldi iteration for computing $M^{1/2}\bm{b}$ in the simulation of particulate suspensions to demonstrate the effectiveness of the proposed method in practical applications.
 Consistently across examples, the Arnoldi approximation is terminated based on the relative residual norm for solving $M\bm{x}=\bm{b}$, with a tolerance of $10^{-2}$.
-In particular, all residuals and relative errors reported in this section are measured in the 2-norm.
-4.1 2D Laplacian
-In this example, we consider the SPD matrix $M$ arising from a five-point stencil discretization of the Laplace equation with homogeneous Dirichlet boundary conditions on the unit square domain.
-The discrete problem is formulated on a uniform numerical grid with grid spacing $h=\frac{1}{n}$, where $n$ denotes the number of divisions along each axis.
-Then, $M$ takes the form of a block-structured matrix:
-$M:=\begin{pmatrix}C&D&0&\cdots&0\\ D&C&D&\cdots&0\\ 0&D&C&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&D&C\end{pmatrix}$
-where, $C$ and $D$ are defined as:
-$C=\operatorname{tridiag}\left(-\frac{1}{h^{2}},\;\frac{4}{h^{2}},\;-\frac{1}{h^{2}}\right)_{(n-1)\times(n-1)},$
-$D=\operatorname{diag}\left(-\frac{1}{h^{2}},-\frac{1}{h^{2}},\dots,-\frac{1}{h^{2}}\right)_{(n-1)\times(n-1)}.$
-In this case, vector $\bm{b}$ is taken to be the vector of all ones.
-The performance of the Arnoldi iteration for computing $M^{1/2}\bm{b}$ is presented in Table 1.
-The numerical experiments confirm that as the condition number of the matrix increases, maintaining a relative error on the order of $10^{-5}$ requires a larger number of iterations.
-The decay of the normalized residual for FOM and the relative error for the Arnoldi method are depicted in Fig.
-2.
-In this example, a normalized residual of $10^{-2}$ for FOM yields a relative error of approximately $10^{-5}$ in the Arnoldi approximation, showing a three order of magnitude improvement.
-This observation indicates that employing the normalized residual of FOM as a stopping criterion is both effective and efficient, enabling accurate results even with a relatively large tolerance.
-To gain further insight into this example, we study the relationship between the relative error, condition number, and iteration count.
-The log-log plot of the relative error decay against the iteration count $k$ for different condition numbers is shown in Fig.
+In particular, all integrals in the bounds are evaluated numerically using integral in MATLAB with a relative tolerance of $10^{-8}$ and an absolute tolerance of $10^{-12}$.
+4.1 Error bounds
+Here, we compare the error bound estimates for cases when the spectrum of the matrix is uniformly distributed versus when the eigenvalues are clustered.
+We test this on both non-Hermitian and Hermitian examples.
+4.1.1 The non-Hermitian case
+Consider a positive definite non-Hermitian matrix $M\in\mathbb{R}^{500\times 500}$, constructed as $M=M_{0}+K$.
+The symmetric component $M_{0}$ is defined using $M_{0}=Q\mathrm{diag}(\Lambda)Q^{T}$, where $Q$ is an orthogonal matrix and $\Lambda$ is a diagonal matrix.
+The matrix $K$ is introduced as a skew-symmetric term derived from a general random matrix $R$:
+$K=\tfrac{1}{2}(R-R^{T}).$
+First, we construct $\Lambda$ such that its eigenvalues are uniformly distributed in the interval $[1,1000]$.
+Fig.
+2 displays the eigenvalues in the complex plane and the distribution of their modulus.
+We see that the eigenvalues of $M$ are widely distributed.
+Figure 2: Eigenvalues of the uniformly-distributed spectrum matrix in the complex plane (left) and their magnitudes (right).
+Setting $\bm{b}$ as an all-ones vector, the error of the Arnoldi process for computing $M^{1/2}\bm{b}$ are shown in Fig.
 3.
-We observe that the error exhibits a decay rate of approximately $\mathcal{O}(k^{-3/2})$ with respect to the iteration count $k$.
-Additionally, we fix $k=25$ and plot the relative error against the condition number on a log-log scale in Fig.
-4.
-As shown, the error decay scales with the condition number $\kappa$ at a rate close to $\mathcal{O}(\kappa^{5/2})$.
-These results are consistent with the conclusions for Hermitian matrices stated in Remark 3.10 and Lemma 3.3 in [1].
-Table 1: Condition number of $M$, iteration counts, and relative error of the Arnoldi iteration for the discrete 2D Laplacian.
-$n$ $\kappa(M)$ Relative Error Iteration Counts $k$
-30 529.97 1.90e-05 29
-40 942.53 1.59e-05 39
-50 1472.96 1.07e-05 50
-60 2121.27 9.98e-06 60
-70 2887.45 7.84e-06 71
-80 3771.51 7.57e-06 81
-90 4773.44 6.31e-06 92
-100 5893.24 6.22e-06 102
-110 7130.92 4.69e-06 114
-Figure 2: Semi-log plot of the normalized residual for FOM (*–) and the relative error for Arnoldi (o–) with different condition number against iteration count $k$ for the discrete 2D Laplacian.
-Figure 3: Log-log plot of relative error decay with varying condition numbers against iteration count $k$ for the discrete 2D Laplacian.
-Figure 4: Log-log plot of relative error with $k=25$ against condition number for the discrete 2D Laplacian.
-4.2 Convection-Diffusion problem
-In this case, $M$ is taken from the discretization of the convection-diffusion problem, which describes transport phenomena in a given domain.
+We observe that the intermediate bound Eq.
+10 remains relatively sharp.
+The overestimation in the a priori bound Eq.
+11 mainly arises from the rough estimate $|\lambda_{i}(H_{k})|\leq\sigma_{\max}(M)$, which is consistent with our earlier discussion in Remark 3.13.
+Figure 3: Error of the Arnoldi process for a non-Hermitian matrix with a uniformly-distributed spectrum.
+Next, let $\Lambda=[\lambda_{\text{cluster}};\lambda_{\text{outlier}}]$, where 95% of the eigenvalues are normally distributed around 1000:
+$\lambda_{\text{cluster}}\sim 1000+100\,\mathcal{N}(0,1).$
+The remaining outliers are drawn from a smaller normal distribution centered near 10:
+$\lambda_{\text{outlier}}\sim 10+5\,\mathcal{N}(0,1).$
+The distribution of eigenvalues and the magnitudes $|\lambda_{i}(M)|$ are displayed in Fig. 4. As shown, most of the $|\lambda_{i}(M)|$ lie around 1000, with only a few small isolated values.
+Figure 4: Eigenvalues of the clustered spectrum matrix in the complex plane (left) and their magnitudes (right).
+The vector $\bm{b}$ is taken to be the average of 100 eigenvectors corresponding to the 100 eigenvalues of largest magnitude.
+Then, the corresponding results of the Arnoldi process are presented in Fig.
+5.
+Due to the clustered eigenvalue distribution, the method exhibits fast convergence, with a priori bounds (11) providing a sharper estimate in this regime as expected.
+Figure 5: Error of the Arnoldi process for a non-Hermitian matrix with a clustered spectrum.
+4.1.2 The Hermitian case
+Next, we construct a $500\times 500$ Hermitian matrix as $M=Q\,\mathrm{diag}(\Lambda)\,Q^{T}$, where the eigenvalues of the diagonal matrix $\Lambda$ follow a uniform distribution over $[1,1000]$.
+In Fig.
+6, we report the error and bounds in Remark 3.14 and Theorem 3.15.
+We observe that, as $k$ increases, the Ritz value $\lambda_{i}(H_{k})$ approaches a true eigenvalue of $M$, leading the bound (9) to become sharper and align more closely with the actual error.
+We also plot the evolution of $\bar{\lambda}_{k+1}(M)$, with $\sigma_{\max}(M)$ shown in Fig.
+7 for comparison.
+It can be observed that when the eigenvalues are uniformly distributed and lie close to one another, bound (13) yields only a minimal improvement over bound (12).
+Figure 6: Error of the Arnoldi process for a Hermitian matrix with a uniformly-distributed spectrum.
+Figure 7: Evolution of $\bar{\lambda}_{k+1}$ for a symmetric matrix with a uniformly-distributed spectrum.
+Next, we take $\Lambda=[\lambda_{\text{cluster}};\,\lambda_{\text{outlier}}]$ with 99% of the eigenvalues tightly clustered near 10, and the remaining 1% sampled as outliers around 1000.
+The same $\bm{b}$ is used.
+The results are displayed in Fig.
+8 and Fig.
+9.
+Clearly, as $k$ increases, the sequence $\bar{\lambda}_{k+1}(M)$ decreases monotonically, and becomes progressively smaller than $\sigma_{\max}(M)$.
+Therefore, as expected in Remark 3.13, when the larger eigenvalues are more widely spread, bound (13) exhibits a tighter estimate compared to bound (12).
+Figure 8: Error of the Arnoldi process for a symmetric matrix with a clustered spectrum.
+Figure 9: Evolution of $\bar{\lambda}_{k+1}$ for a symmetric matrix with a clustered spectrum.
+4.2 Scaling behavior
+In this example, we test the algorithm’s dependence on the number of Arnoldi iterations $k$ and the largest singular value $\sigma_{\max}(M)$.
+Here, we construct $M$ from the discretization of the convection-diffusion problem.
 The continuous convection-diffusion operator is defined as
-$Lu:=\eta u^{{\prime}{\prime}}(x)+u^{\prime}(x),\quad x\in[0,1],$
+$Lu:=-\eta u^{{\prime}{\prime}}(x)+u^{\prime}(x),\quad x\in[0,1],$
 subject to homogeneous Dirichlet boundary conditions.
@@
 $\eta\frac{u_{i+1}-2u_{i}+u_{i-1}}{h^{2}}+\frac{u_{i}-u_{i-1}}{h},$
-where $u_{i}$ represents the numerical approximation to $u(x_{i})$.
-Let $M$ be the tridiagonal matrix obtained from the above finite-difference discretization of $L$.
-We take $\bm{b}$ to be the vector with all entries equal to one, and fix $\eta$ = 0.1.
-The results for approximating $M^{1/2}\bm{b}$ are summarized in Table 2 and Fig.
-5, and the general trends align with the conclusions drawn form the 2D Laplacian example.
-However, the matrix in this case is no longer symmetric.
-From the plot of the error versus iteration count in Fig.
-6, we see that as the condition number increases, the convergence rate closely follows a $k^{-3/4}$ relationship with respect to the iteration count.
+where $u_{i}$ represents the numerical approximation to $u(x_{i})$. Let $M$ be the tridiagonal matrix obtained from the above finite-difference discretization of $L$, which is non-Hermitian.
+Here, we take $\bm{b}$ to be the vector with all entries equal to one, and choose $\eta$ = 0.1.
+In this case, $M$ exhibits a relatively dispersed eigenvalue distribution.
+We compare different error bounds for approximating the square root of $M$ using the Arnoldi process described above (see Fig.
+10).
+Similar phenomena as in Section 4.1 can also be seen here.
+Although the bound Eq.
+11 is loose, its overall trend matches that of the true error, providing insights into the inherent scaling behavior.
+Figure 10: Error of the Arnoldi process for a convection-diffusion problem with $n=500$.
+Next, we examine how the error depends on $\sigma_{\max}(M)$ and $k$, and whether the observed behavior follows the scaling predicted by bound (11).
+To this end, we introduce the scaling term derived from (11), defined as
+$\frac{2^{1/4}\pi}{\Gamma\!(3/4)}\frac{2k-3}{2k}\frac{\left\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(f;M,\bm{b})\right\|}{\left\|\bm{\xi}_{0}^{k}\right\|}.$
+In the following tests, we use the stopping criterion in Remark 3.10 with tol = 0.05.
+The results of approximating $M^{1/2}\bm{b}$ for matrices of varying dimensions arising from the convection-diffusion problem are summarized in Table 1.
+We see that as $n$ increases, maintaining the same accuracy requires a larger number of iterations.
+The log-log plot of the scaling term decaying against the iteration count $k$ is shown in Fig.
+11.
+The convergence rate exhibits a decay rate of approximately $\mathcal{O}(k^{-3/4})$ with respect to the iteration count $k$.
 This observation is consistent with the theoretical result established in Theorem 3.11.
-Meanwhile, for a fixed $k=480$, Fig.
-6 indicates that for large condition numbers, the observed trend closely follows $\kappa^{7/2}$.
-Table 2: Condition number of $M$, iteration counts, and relative error of the Arnoldi iteration for the convection-diffusion problem.
-$n$ $\kappa(M)$ Relative Error Iteration Counts $k$
-500 67510.90 3.99e-08 495
-600 97087.72 3.31e-08 594
-700 132022.80 2.85e-08 693
-800 172316.14 2.53e-08 792
-900 217967.73 2.29e-08 891
-1000 268977.57 2.10e-08 990
-1100 325345.68 1.94e-08 1089
-1200 387072.03 1.81e-08 1188
-1300 454156.65 1.71e-08 1287
-Figure 5: Semi-log plot of the normalized residual for FOM (*–) and the relative error for Arnoldi (o–) with different condition number against iteration count $k$ for the convection-diffusion problem.
-Figure 6: Log-log plot of relative error decay with varying condition numbers against iteration count $k$ for the convection-diffusion problem.
-Figure 7: Log-log plot of relative error with $k=480$ against condition number for the convection-diffusion problem.
-4.3 General matrices
-Next, we consider positive definite but non-symmetric dense matrices. The matrix $M$ is constructed as follows:
-(18) $M=M_{0}+\alpha I=USV^{T}+\alpha I,$
-where $M_{0}$ is a low-rank matrix with rank $m$, $U$ and $V$ are orthonormal matrices generated randomly, and $S$ is an $m\times m$ diagonal matrix whose entries decay logarithmically from 1 to $\frac{1}{\beta}$, where $\beta$ is a given constant.
-A shift term, $\alpha I$, is introduced to ensure positive-definiteness of $M$.
-Specifically, the shift parameter $\alpha$ is defined as: $\alpha=-(1+5\times 10^{-3})\min\operatorname{Re}(\lambda_{\min}(M_{0})).$ As a result, the condition number of the $M$ is influenced by both $\alpha$ and $\beta$.
-In this test, we set the rank to $m=500$, and the size of the matrix $M$ is $5000\times 5000$.
-The parameter $\beta$ takes values of 400, 600, 800, 1000, 1200, 1400, 1600, 1800, 2000.
-As before, $\bm{b}$ is taken to be the vector of all ones.
-In Table 3, we display the corresponding results.
-As the relative error remains around $10^{-5}$, the number of iterations generally increases with the condition number.
-Similarly, the normalized residual for solving $M\bm{x}=\bm{b}$ provides an effective stopping criterion for the Arnoldi iteration in the evaluation of $M^{1/2}\bm{b}$ (Fig.
-8).
-As seen in Fig.
-9, the error decay of the Arnoldi method closely follows the reference slope before the transition to superlinear convergence.
-However, we observe that with the iteration count fixed at $k=35$, the fitted slope is approximately $0.9$, which is much smaller than the theoretical value of $7/2$.
-These observations indicate that the theoretical results established in Theorem 3.13 yield a reasonably tight estimate for the iteration count $k$, but only a loose upper bound with respect to the condition number for certain type of matrices.
-This may be due to the fact that the bounds could depend not only on the condition number and iteration counts but also on intrinsic spectral characteristics or a specific structure of the matrices.
-Table 3: Condition number of $M$, iteration counts, and relative error of the Arnoldi iteration for general matrices constructed via Eq. 18.
-$\kappa(M)$ Relative Error Iteration Counts $k$
-2192.38 2.40e-05 38
-5375.35 2.90e-05 41
-9325.42 2.62e-05 56
-28826.11 3.86e-05 38
-32437.66 2.45e-05 47
-47395.74 2.87e-05 58
-52845.93 6.31e-05 36
-166131.16 5.27e-05 63
-188372.52 1.03e-04 51
-Figure 8: Semi-log plot of the normalized residual for FOM (*–) and the relative error for Arnoldi (o–) with different condition number against iteration count $k$ for general matrices constructed via Eq.
-18.
-Figure 9: Log-log plot of relative error decay with varying condition numbers against iteration count $k$ for general matrices constructed via Eq. 18.
-Figure 10: Log-log plot of relative error with $k=35$ against condition number for general matrices constructed via Eq. 18.
+Additionally, we fix $k=300,500,700$ and plot the scaling term against the largest singular value on a log-log scale in Fig.
+12.
+The results show that the fitted slopes vary with $k$, but, as expected, they all fall below the theoretical prediction.
+Table 1: Condition number of $M$ and iteration counts $n$ and errors of the Arnoldi process for a convection-diffusion problem.
+$n$ $\sigma_{\max}(M)$ error Iteration Counts $k$
+1000 202320.64 0.03083 889
+1200 291138.58 0.03053 1071
+1400 396074.49 0.03061 1253
+1600 517128.36 0.03090 1435
+1800 654300.20 0.03132 1617
+2000 807590.00 0.03132 1800
+Figure 11: Log-log plot of error decay with varying condition numbers against iteration count $k$ for a convection-diffusion problem.
+Figure 12: Log-log plot of the error with $k=150$ against the largest singular value for a convection-diffusion problem.
 4.3 Practical application
-The study of particulate suspensions is of great interest, and plays critical roles in diverse fields such as materials design [27], biological systems [27], micro-robotics [37], and food processing [21].
+The study of particulate suspensions is of great interest, and plays critical roles in diverse fields such as materials design [makey2020universality], biological systems [makey2020universality], micro-robotics [yang2020reconfigurable], and food processing [kim2020modeling].
 Immersed in a fluid environment, particles’ mobility is dictated by the hydrodynamic interactions between the particles.
 At micro scales, particles are also subject to thermal fluctuations, in the form of Brownian motion.
-The configuration of particles evolves over time and is governed by the overdamped limit of Langevin dynamics [14].
+The configuration of particles evolves over time and is governed by the overdamped limit of Langevin dynamics [ermak1978brownian].
 Thus, the computation of particles’ Brownian displacements crucially rely on the mobility matrix, and its square root, both dependent on all particles’ instantaneous positions.
 Efficiently computing these two matrices remains a significant computational challenge.
-Since hydrodynamic interactions depend on the separation distance between particles, the particle interactions weaken as the distance increases [25].
+Since hydrodynamic interactions depend on the separation distance between particles, the particle interactions weaken as the distance increases [ma2022fast].
 As a result, far-field blocks in the mobility matrix exhibit low-rank structure, and can then be efficiently approximated.
-To this end, the Hydrodynamic Interaction Graph Neural Network (HIGNN) framework, proposed in [25, 24], maps the many-body hydrodynamic interactions onto a graph neural network.
-By employing hierarchical-matrices using adaptive cross approximation [17], the computational complexity for using HIGNN to compute particles’ velocities given any applied forces can be reduced to nearly linear [26].
+To this end, the Hydrodynamic Interaction Graph Neural Network (HIGNN) framework, proposed in [ma2022fast, ma2024shape], maps the many-body hydrodynamic interactions onto a graph neural network.
+By employing hierarchical-matrices via adaptive cross approximation [hackbusch2015hierarchical], the computational complexity for using HIGNN to compute particles’ velocities given any forces can be reduced to nearly linear [ma2025mathcal].
 In the following tests, $M$ is the mobility matrix obtained from the HIGNN model and has been preprocessed into a hierarchical-matrix form.
@@
 The number of particles, $N$, considered here varies from $20^{3}$ to $100^{3}$.
-For such large systems, direct computation of the square root of the mobility matrix $M$ can be computationally expensive.
-To address this and evaluate the accuracy of the Arnoldi iteration in this context, we apply the Arnoldi process twice to approximate $M\bm{b}=M^{1/2}(M^{1/2}\bm{b})$.
+For such large systems, direct computation of the square root of the mobility matrix $M$ can be computationally prohibitive.
+To address this challenge and also validate the accuracy of the Arnoldi approximation in this practical setting, we apply the Arnoldi process twice to approximate $M\bm{b}=M^{1/2}(M^{1/2}\bm{b})$.
 Thus, the relative error is defined as $\frac{\|M\bm{b}-\operatorname{Arn}_{k}(x^{1/2};M,\operatorname{Arn}_{k}(x^{1/2};M,\bm{b}))\|}{\|M\bm{b}\|}$.
-Figure 11: Snapshots of 8000 particles sedimenting in an unbounded domain at times $t=0.005$, $0.25$, and $1$, respectively, simulated using the HIGNN.
+Figure 13: Snapshots of 8000 particles sedimenting in an unbounded domain at times $t=0.005$, $0.25$, and $1$, respectively, simulated using the HIGNN [ma2025mathcal].
 The color is correlated to the particles’ velocity magnitude.
-Table 4: Arnoldi Approximation for lattice structure system with constant particle concentration.
-$N$ Iteration Counts $k$ Relative Error
+Table 2: Iteration counts and errors of the Arnoldi process for a particulate suspension problem with varying particle number NN at fixed concentration.
+NN Iteration Counts $k$ Relative Error
 $20^{3}$ 10 2.06699e-04
@@
 The snapshots at the 1st, 100th, and 200th time steps from the simulation of a $20^{3}$ particle system are presented in Fig.
-11.
+13.
 Letting $M$ be the mobility matrix at the 100th time step, Table 2 records the number of iterations $k$ required to compute $M^{1/2}\bm{b}$ and the relative error as the number of particles $N$ increases from $20^{3}$ to $100^{3}$.
-Note that in this scenario, since the initial spacing between particles is kept the same as $N$ increases, a consistent particle concentration is maintained.
+Note that in this scenario, since the initial spacing between particles is kept the same as $N$ increases, a fixed particle concentration is maintained.
 One sees that as the number of particles increases, the number of iterations required to compute the matrix square root applied to a vector remains relatively low and does not grow significantly, while maintaining a relative error on the order of $10^{-4}$.
@@
 Furthermore, we depict the relationship between the number of particles and the runtime on a log-log scale in Fig.
-12.
+14.
 The plot implies that the Arnoldi approximation, based on the hierarchical matrix representation, exhibits an $O(N\log N)$ time complexity.
-Figure 12: Log-log plot of runtime against number of particles for lattice structure system.
+Figure 14: Log-log plot of runtime versus number of particles NN at fixed particle concentration.
 Next, we consider $40^{3}$ particles arranged in a similar initial cubic lattice as above, but with the grid spacing $d$ varying from 3 to 5 (still perturbed by a displacement of 0.2).
-With this change, we consider different particle concentrations; when $d$ increases, the particle concentration decreases, thus lowering the required rank of the approximation to the corresponding blocks.
+With this change, we can examine varying particle concentrations; as $d$ increases, the particle concentration decreases, thus lowering the required rank of the approximation to the corresponding blocks.
 As shown in Table 3, lower particle concentration leads to reduced runtime for the Arnoldi approximation.
-However, the required number of iterations remains relatively small and stable compared to the matrix size.
+Also, the required number of iterations remains relatively small and stable compared to the matrix size.
 In this test, $M$ corresponds to the mobility matrix attained at the 50th time step.
-Table 5: Arnoldi Approximation for lattice structure with different particle concentration.
+Table 3: Iteration counts, errors, and runtime of the Arnoldi process for a particulate suspension problem with varying particle concentration at $N=40^{3}$.
 $d$ Iteration Counts $k$ relative error runtime
-$3$ 12 1.93793e-04 362.4649
+33 12 1.93793e-04 362.4649
 $3.5$ 12 1.81240e-04 386.0504
-$4$ 10 1.88017e-04 319.9872
+44 10 1.88017e-04 319.9872
 $4.5$ 10 1.86118e-04 307.9138
-$5$ 10 1.83476e-04 298.9904
+55 10 1.83476e-04 298.9904
 5 Concluding remarks
-In this work, we derive a priori error estimates of the Arnoldi process for approximating the matrix square root applied to a vector, extending the results in [6] for Hermitian matrices to non-Hermitian cases.
-Moreover, we provide an error bound based on a data-sparse representation of the matrix.
-Through numerical experiments, we investigate the convergence behavior of the Arnoldi iteration for computing $M^{1/2}\bm{b}$ across different matrix types.
-The observations indicate that the theoretical results presented in Theorem 3.8 appear to be sharp in terms of iteration count, yet remain a loose upper bound with respect to the condition number for certain classes of matrices.
-Additionally, we consider the mobility matrix arising from particulate suspension simulations, where the matrix is approximated into a hierarchical matrix.
+In this work, we derive error estimates of the Arnoldi process for approximating the matrix square root applied to a vector, extending the results reported in [chen2022error] for Hermitian matrices to non-Hermitian cases.
+In addition, we also provide a sharper bound for the Hermitian case, improving the state-of-the-art result.
+Moreover, we provide an error bound based on a perturbed matrix.
+Through a series of numerical experiments, we systematically assess the convergence behaviors of the Arnoldi iteration for computing $M^{1/2}\bm{b}$ across different matrix types.
+The findings indicate that the theoretical results presented in Theorem 3.8 appear to be sharp in terms of iteration count, yet remain a loose upper bound with respect to the largest singular value for certain classes of matrices.
+Furthermore, we consider the mobility matrix arising from particulate suspension simulations, where the matrix is approximated into a hierarchical form.
 The numerical results demonstrate the effectiveness of our approach for large-scale problems.
-Future extensions of this paper include deriving a sharper dependence on the condition number for approximating the matrix square root applied to a vector via Arnoldi iteration, as well as investigating the error behavior of GMRES in the approximation of $f(M)\bm{b}$ for general functions $f$.
+Future extensions of this work include deriving a sharper dependence on the largest singular value for approximating the matrix square root applied to a vector via Arnoldi iteration, as well as investigating the error behaviors of GMRES in the approximation of $f(M)\bm{b}$ for general functions $f$.
 Acknowledgments
 The authors would like to thank Zhan Ma and Zisheng Ye for their valuable discussions and assistance with the code implementation.
-The authors also gratefully acknowledge support from the National Science Foundation under Grant No.
+The authors also gratefully acknowledge funding support from the National Science Foundation under Grant No.
 DMS-2208267 and the Army Research Office under Grant No.